### v1

# A YTD correspondence for constant scalar curvature metrics

Abstract
Given a compact Kähler manifold, we provide a uniform YTD correspondence that characterizes the existence of a unique constant scalar curvature metric using ample/Kähler test configurations. The main point of our approach is to replace the $K$ energy with much better behaving variants, whose properties are similar to the Ding energy from the Fano case. Connections with the non-Archimedean point of view on K-stability are also briefly discussed.
Contents
1 The main results
Let $(X,\omega)$ be a compact Kähler manifold. In the polarized setting, classical formulations of the Yau–Tian–Donaldson (YTD) conjecture aim to characterize the existence of constant scalar curvature Kähler (cscK) metrics using invariants associated to test configurations [93, 85, 51]. Our work fits within this general framework. While we do not address any specific version of the YTD conjecture from the literature, we do establish a YTD correspondence: we characterize existence of cscK metrics, using ample test configurations and a novel algebraic invariant. Our approach works equally well in the transcendental setting.
To start, we recall that there is a one-to-one correspondence (up to a constant) between Kähler metrics cohomologous to $\omega$ and the space of Kähler potentials:
$\mathcal{H}_{\omega}:=\{u\in C^{\infty}(X)\textup{ s.t. }\omega_{u}:=\omega+\mathrm{dd^{c}}u>0\},$
where $\mathrm{dd^{c}}:=\frac{i}{2\pi}\partial\bar{\partial}.$ The study of canonical representatives on $\mathcal{H}_{\omega}$ goes back to questions of Calabi [24] and the famous theorem by Yau [92]. Relevant to our work, the cscK equation can be formulated as follows:
$S_{\omega_{u}}=\overline{S},\ \ \ u\in\mathcal{H}_{\omega},$ (1)
where $S_{\omega_{u}}$ is the scalar curvature of $\omega_{u}$, $\overline{S}=\frac{n}{V}\int_{X}\operatorname{Ric}\omega\wedge\omega^{n-1}$ is the average scalar curvature, and $V=\int_{X}\omega_{u}^{n}$ is the total Kähler volume.
Solutions to (1) are critical points of Mabuchi’s K energy functional $K:\mathcal{H}_{\omega}\to\mathbb{R}$ [64]:
$K(u)={Ent}(u)-n{I}_{\operatorname{Ric}(\omega)}(u)+\bar{S}{I}(u),$ (2)
where ${Ent}(u)=\frac{1}{V}\int_{X}\log(\omega^{n}_{u}/\omega^{n})\omega^{n}_{u}$ is the entropy of $\omega^{n}_{u}$ with respect to $\omega^{n}$, whereas ${I},{I}_{\textup{Ric}(\omega)}:\mathcal{H}_{\omega}\to\mathbb{R}$ are the Monge-Ampère (Aubin-Yau) energy and its Ricci-twisted version:
${I}(u)=\frac{1}{(n+1)V}\sum_{j=0}^{n}\int_{X}u\omega^{j}_{u}\wedge\omega^{n-j},\ \ \ {I}_{\textup{Ric}(\omega)}(u)=\frac{1}{nV}\sum_{j=0}^{n-1}\int_{X}u\textup{Ric}(\omega)\wedge\omega_{u}^{j}\wedge\omega^{n-j-1}.$ (3)
$K^{\beta}$ energy and transcendental quantization.
Traditionally, the most difficult term in (2) is the entropy which is only lower semi-continuous (lsc) with respect to the $d_{1}$ metric of $\mathcal{H}_{\omega}$ (introduced in [32]). To deal with this, we propose to replace the entropy with a damper version, having much better properties. For $\beta>0$ we introduce the $\beta$-entropy:
$Ent^{\beta}(u):=\sup_{v\in\mathcal{H}_{\omega}}\left(-\log\int_{X}e^{\beta(v-u)}\frac{\omega^{n}}{V}+\beta(I(v)-I(u))\right).$ (4)
Recalling the formula $Ent(u)=\sup_{w\in C^{0}(X)}\left(-\log\frac{1}{V}\int_{X}e^{w}{\omega^{n}}+\frac{1}{V}\int_{X}w\omega_{u}^{n}\right)$ ([8, Proposition 2.10]) and the basic inequality $I(v)-I(u)\leq\frac{1}{V}\int_{X}(v-u)\omega_{u}^{n}$, we immediately see that $Ent^{\beta}(u)\leq Ent(u)$. In Corollary 3.11 we will see that $Ent^{\beta}(u)\nearrow Ent(u)$ as $\beta\nearrow\infty$.
It may seem surprising at first, but our formulation of the $\beta$-entropy is intimately tied to Berman’s speculations from [6, §3.1], where he suggests to view solutions $u^{\beta}$ to the following Aubin–Yau type Monge–Ampère equation as ‘transcendental quantizations’ of $u\in\mathcal{H}_{\omega}$, with $\beta^{-1}$ serving as the Planck constant:
$(\omega+\mathrm{dd^{c}}u^{\beta})^{n}=e^{\beta(u^{\beta}-u)}\omega^{n}.$
Indeed, one can ‘linearize’ this equation to obtain a formal asymptotic expansion for $u^{\beta}$:
$u^{\beta}\sim u+\frac{1}{\beta}\log\frac{\omega^{n}_{u}}{\omega^{n}}+\frac{1}{\beta^{2}}(\mathrm{tr}_{\omega_{u}}\operatorname{Ric}(\omega)-S(\omega_{u}))+O\Big(\frac{1}{\beta^{3}}\Big),\ \beta\to\infty.$ (5)
Though not used (nor proved) in this work, this expansion gives a transcendental analogue of the classical Bergman kernel asymptotic expansions going back to Tian [83]. Intriguingly, the scalar curvature appears in the above expansion, as in the Bergman kernel expansion. For classical Bergman kernels, this point was crucially used by Donaldson [50, 52], showing that cscK metrics are closely related to the Chow stability in algebraic geometry.
For our $\beta$-entropy, the following crucial formula is proved in Proposition 3.6:
$Ent^{\beta}(u)=\beta(I(u^{\beta})-I(u)).$ (6)
The expression on the right hand side is reminiscent of Donaldson’s $\mathcal{L}$-function from [52] in the classical Bergman kernel setting. Because of this analogy, we call $Ent^{\beta}$ the (transcendentally) quantized entropy at level $\beta$. The corresponding quantized $K^{\beta}$ energy functional $K^{\beta}:\mathcal{H}_{\omega}\to\mathbb{R}$—satisfying $K^{\beta}(u)\nearrow K(u)$ as $\beta\nearrow\infty$—is defined as follows:
$K^{\beta}(u)={Ent}^{\beta}(u)-n{I}_{\textup{Ric }\omega}(u)+n\bar{S}{I}(u).$ (7)
Due to Chen-Cheng, existence of cscK metrics is equivalent to properness of the K energy [26, 27], which we show to be further equivalent with properness of the ‘quantized’ $K^{\beta}$ functionals for $\beta$ big enough (Theorem 4.4). This is in line with quantization philosophy extensively explored in the algebraic setting (see e.g. [83, 84, 50, 52, 69, 70, 78, 96]). Our argument relies on estimates reminiscent of the partial $C^{0}$ estimate (Proposition 4.1), originally introduced by Tian [84] in the Fano setting.
In addition, we show that the $K^{\beta}$ energy and its radial version are continuous with respect to the $d_{1}$ and $d_{1}^{c}$ metrics, respectively (Lemma 3.9, Theorem 5.2). The analogous properties are known to hold for the Ding energy ([39, Proposition 5.19], [41, Theorem 1.4] [21, Proposition 3.15]) but fail for the $K$ energy. Thus, while the properness profile of $K^{\beta}$ agrees with that of the $K$ energy, its crucial continuity profile is in line with the much better behaving Ding energy known from the Fano case.
Since, in the Fano case, the favorable properties of the Ding energy have driven major advances in $K$-stability, moduli theory, $\delta$-invariants, and other areas [14, 28, 86, 60, 53, 54, 16, 91, 63, 96, 42], we are motivated to make extensive use of $K^{\beta}$ functionals in this work.
A YTD correspondence.
In the case of an ample line bundle $L\to X$ with hermitian metric $h$ whose curvature equals our Kähler metric $\omega$, to test K-stability, one considers test configurations. Historically this notion is due to Ding–Tian and Donaldson and involves one parameter degeneration of the structure $(X,L)$ [49, 85, 51]. We give the precise definition of ample test configurations in §2. One may think of these as finitely generated $\mathbb{Z}$-filtrations of the section ring $R(X,L^{k})$ for some $k\in\mathbb{N}$. This perspective was first considered in [88, 80] and it was shown in [17, Proposition 2.15] that it is equivalent with the more classical notion introduced in [85, 51].
To an ample test configuration $\mathcal{T}$, due to work initiated by Phong-Sturm, it is possible to associate a $C^{1,1}$ Mabuchi geodesic ray $0,\infty)\ni t\to u^{\mathcal{T}}_{t}\in\mathcal{H}_{\omega}^{1,1} [$72, 73, 76, 5, 30].
We introduce invariants associated with $K^{\beta}$ and $\mathcal{T}$, using asymptotic slopes along $\{u^{\mathcal{T}}_{t}\}$:
[…]
Here $J(u)=\frac{1}{V}\int_{X}u\omega^{n}-I(u)$ is the well known $J$ functional, that is known to be convex.
Given that $\mathcal{T}$ is algebraic, it is natural to seek an interpretation of the associated invariants purely in algebraic terms. For instance, $J_{\mathcal{T}}$ admits such a description via certain intersection numbers attached to $\mathcal{T}$ [71, 17]. For $K^{\beta}_{\mathcal{T}}$, an explicit intersection-theoretic formula will be provided in a sequel [43] (see formula (50)). In the present work we provide a more readily available divisorial threshold–type interpretation, in the spirit of delta and alpha invariants (see Corollary 8.2), already emphasizing the algebraic nature of $K^{\beta}_{\mathcal{T}}$.
Given $\beta>0$, we say that $(X,L)$ is uniformly $K^{\beta}$-stable if there exists $\gamma>0$ such that
$K^{\beta}_{\mathcal{T}}\geq\gamma J_{\mathcal{T}},$ (9)
for all ample test configurations $\mathcal{T}$.
The first main result of this paper is the following uniform YTD correspondence for cscK metrics. To our knowledge this is the first characterizing criterion with ample test configurations as test objects:
Theorem 1.1.
Let $(X,L)$ be a polarized Kähler manifold. The following are equivalent.
(i) $(X,L)$ admits a unique cscK metric in $c_{1}(L)$.
(ii) $(X,L)$ is uniformly $K^{\beta}$-stable for some $\beta>0$.
Given the nature of our techniques, our argument also yields a YTD correspondence for transcendental Kähler manifolds $(X,\omega)$ with virtually identical statement (see Theorem 6.2).
To explain the proof of Theorem 1.1, let us recall a few concepts. By $(\mathcal{E}^{1},d_{1})$ we denote the $d_{1}$ metric completion of $\mathcal{H}_{\omega}$, explored in [32], with the strongly topology of the space $\mathcal{E}^{1}$ already investigated in [8]. The space $(\mathcal{E}^{1},d_{1})$ is a geodesic metric space [32, Theorem 2]. Our functionals $K$, $K^{\beta}$, and $J$ extend naturally to $\mathcal{E}^{1}$, preserving their $d_{1}$-(semi)continuity properties [7, 11]. Moreover, the extended $K^{\beta}$ continues to satisfy the formulas (4) and (6).
By $\mathcal{R}^{1}$ we denote the space of geodesic rays $t\mapsto u_{t}$ inside $\mathcal{E}^{1}$, that emanate from the background potential $0\in\mathcal{H}_{\omega}$ (notation: $\{u_{t}\}\in\mathcal{R}^{1}$). Due to Buseman convexity of $d_{1}$ ([11, Proposition 5.1]) one can introduce the chordal distance for $\{u_{t}\},\{v_{t}\}\in\mathcal{R}^{1}$ [37, 27]:
$d_{1}^{c}(\{u_{t}\},\{v_{t}\})=\lim_{t\to\infty}\frac{d_{1}(u_{t},v_{t})}{t}.$ (10)
Thus defined, $(\mathcal{R}^{1},d_{1}^{c})$ is a complete geodesic metric space [37, Theorem 1.4].
An important subspace of $(\mathcal{R}^{1},d_{1}^{c})$ is $\mathcal{R}^{1}_{\mathcal{I}}$, defined as the $d_{1}^{c}$-closure of rays $\{u_{t}^{\mathcal{T}}\}$ induced by test configurations $\mathcal{T}$. In [10] the authors refer to such rays simply as maximal. This space admits a natural identification with $\mathcal{E}^{1,NA}$ (see [40, Theorem 1.2] and the discussion preceding it). For more detailed treatments of this topic, we refer to [37, 10, 40, 34].
Given $\{u_{t}\}\in\mathcal{R}^{1}$ it is convenient to introduce the radial version of these functionals:
$K\{u_{t}\}=\lim_{t\to\infty}\frac{K(u_{t})}{t},\ \ \ K^{\beta}\{u_{t}\}=\liminf_{t\to\infty}\frac{K^{\beta}(u_{t})}{t},\ \ \ J\{u_{t}\}=\lim_{t\to\infty}\frac{J(u_{t})}{t}.$
In a recent exciting development, C. Li showed that rays $\{u_{t}\}\in\mathcal{R}^{1}$ satisfying $K\{u_{t}\}<\infty$ must lie in $\mathcal{R}^{1}_{\mathcal{I}}$ [61, Theorem 1.2].
The proof of Theorem 1.1 relies on four central components. From the existing literature, we will need that existence cscK metrics is equivalent with properness/geodesic stability of the K energy [26, 27], together with the aforementioned result of C. Li [61]. The remaining two components are facts we establish in this work: that the properness of the K energy is equivalent to the properness of the $K^{\beta}$ energy for sufficiently large $\beta$ (Theorem 4.4); and that the radial $K^{\beta}$ energy is $d_{1}^{c}$-continuous (Theorem 5.2).
Once the key ingredients are in place, the argument unfolds naturally. We present it here, as this simple proof serves as the skeleton of the paper (cf. [10, 53] for related ideas in the Fano case).
Proof of Theorem 1.1.
Assume that (i) holds. By [12, Theorem 1.5] the K energy is proper, i.e., $K(u)\geq\gamma J(u)-\beta,\ u\in\mathcal{E}^{1}$ for some $\gamma,\beta>0$. By Theorem 4.4 we obtain that $K^{\beta}$ is proper too, for $\beta>0$ big enough. By our definition of $K^{\beta}_{\mathcal{T}}$ and $J_{\mathcal{T}}$, (ii) immediately follows.
Now we assume that (ii) holds. By [61, Theorem 1.2] any ray with bounded K energy slope is in $\mathcal{R}^{1}_{\mathcal{I}}$, i.e., it is approximable by test configurations. By [27] (cf. [37, Theorem 1.9]), it suffices to show existence of $\gamma>0$ such that
$\mathcal{K}\{u_{t}\}\geq\gamma J\{u_{t}\},\ \ \ \{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}.$
Let $\{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}$ and $\{u^{\mathcal{T}_{j}}_{t}\}$ be $d_{1}^{c}$-approximating rays of ample test configurations $\mathcal{T}_{j}$.
Condition (ii) implies that for some $\beta>0$ and $\gamma>0$ we have that
$K^{\beta}\{u^{\mathcal{T}_{j}}_{t}\}=K^{\beta}_{\mathcal{T}_{j}}\geq\gamma J_{\mathcal{T}_{j}}=\gamma J\{u^{\mathcal{T}_{j}}_{t}\}.$
The functional $J\{\cdot\}$ is known to be $d_{1}^{c}$-continuous (as follows from [32, Lemma 4.15] and [10, Lemma A.2]). $K^{\beta}\{\cdot\}$ also has this property by Theorem 5.2. Thus, using $K\geq K^{\beta}$ (see (19)), we can let $j\to\infty$ in the above inequality to obtain $K\{u_{t}\}\geq K^{\beta}\{u_{t}\}\geq\gamma J\{u_{t}\}.$
Given that Theorem 1.1 introduces a novel stability condition in the context of YTD type results, the second goal of the paper is to explore the meaning of $K^{\beta}$-stability, and its relationship to more classical notions of stability previously considered in the literature.
Radial $K^{\beta}$ and algebraic interpretation of $K^{\beta}_{\mathcal{T}}$.
As noted above, $K^{\beta}(u)\nearrow K(u)$ as $\beta\nearrow\infty$ (see (19)). In our next main result we show that a similar quantization property holds for the radial version of these functionals as well:
Theorem 1.2 (=Theorem 7.3).
Let $\{u_{t}\}\in\mathcal{R}^{1}$. Then
$\lim_{\beta\to\infty}K^{\beta}\{u_{t}\}=K\{u_{t}\}.$
Given an ample test configuration $\mathcal{T}$, recall that $K_{\mathcal{T}}=K\{u^{\mathcal{T}}_{t}\}$ agrees with the non-Archimedean K energy, which also equals the Futaki invariant of $\mathcal{T}$ if the central fiber of $\mathcal{T}$ is reduced [17, (0.1)]. By the above result, the invariants $K^{\beta}_{\mathcal{T}}$ from (8) converge to this well-known algebro-geometric invariant:
Corollary 1.3.
Let $(\mathcal{X},\mathcal{L})$ be an ample test configuration for $(X,L)$. Then
$\lim_{\beta\to\infty}K^{\beta}_{\mathcal{T}}=K_{\mathcal{T}}.$
We further show in Corollary 8.2 that $K^{\beta}_{\mathcal{T}}$ can be computed algebraically, in terms of log canonical thresholds and intersection numbers associated to $\mathcal{T}$, showing similarities with Berman’s algebraic formula for the slope of the Ding energy along test configurations [5, (3.19)]. At the end of §8 we also establish that, in the Fano case, uniform $K^{\beta}$-stability is equivalent to uniform Ding stability. So in light of Theorem 1.1, our approach also gives a new proof for the uniform YTD conjecture for Fano manifolds. In general, the availability of a threshold-type expressions for $K^{\beta}_{\mathcal{T}}$ suggests promising directions for investigating $K^{\beta}$-stability through algebro-geometric techniques, with the goal of developing novel existence criterions for cscK metrics, in line with recent advances in the Fano setting [60, 53, 54, 16, 1, 3, 90].
Our aforementioned results all rely on a general formula for $Ent^{\beta}\{u_{t}\}$ on bounded geodesic rays, which is a radial version of the identity (4) (Theorem 7.1):
$Ent^{\beta}\{u_{t}\}=\sup_{\{w_{t}\}}\left(L^{\beta}(\{w_{t}\},\{u_{t}\})+\beta(I\{w_{t}\}-I\{u_{t}\})\right),$ (11)
where the sup is over geodesic rays $\{w_{t}\}$ arising from ample test configurations. Here $L^{\beta}(\{w_{t}\},\{u_{t}\})$ is the following generalization of the radial Ding functional:
$L^{\beta}(\{v_{t}\},\{u_{t}\}):=\liminf_{t\to\infty}\left(-\frac{1}{t}\log\int_{X}e^{\beta(v_{t}-u_{t})}\frac{\omega^{n}}{V}\right),$
which can be further expressed as an integrability threshold (see Proposition 7.2).
Non-Archimedean interpretation of $Ent^{\beta}$.
When dealing with non-Archimedean data, it is possible to describe the radial $Ent^{\beta}$ functional using a concise formula. For details on terminology and non-Archimedean geometry we refer to [17] and §8.
Let $u\in\mathcal{H}^{NA}$ be a non-Archimedean metric in the spirit of [17], encoding the data of ample test configurations using divisorial information. As such, to $u$ one can again associate a $C^{1,1}$ geodesic ray $\{u_{t}\}$. As a direct consequence of (11), we obtain the following formula for the radial $\beta$-entropy of $\{u_{t}\}$ (Corollary 9.2):
$Ent^{\beta}\{u_{t}\}=\sup_{w\in\mathcal{H}^{NA}}\left(L^{NA}(\beta(u-w))+\beta(I^{NA}(w)-I^{NA}(u))\right),$ (12)
where $L^{NA}$ and $I^{NA}$ are non-Archimedean functions defined in [20, Section 2].
Since it is purely a non-Archimedean expression, one may define the right hand side of (12) as $Ent^{\beta,NA}(u)$, the non-Archimedean $\beta$-entropy.
It will easily follow from our analysis that the above formula also holds for $u\in\mathcal{H}^{NA}_{\mathcal{M}}$, the non-Archimedean potentials arising from a model (filtration) of C. Li [61, Definition 2.7]. More importantly, using [22, (4.5)] we immediately obtain the inequality $Ent^{\beta}\{u_{t}\}\leq Ent^{NA}(u),\ u\in\mathcal{H}^{NA}_{\mathcal{M}}.$ Now we can use [61, Proposition 6.3] and Theorem 5.2 to conclude that, more generally:
$Ent^{\beta}\{u_{t}\}\leq Ent^{NA}(u),\ \ u\in\mathcal{E}^{1,NA}.$
Adding the terms coming from the radial Monge-Ampère energy and its twisted version, we immediately obtain that $K^{\beta}\{u_{t}\}\leq K^{NA}(u),\ \ u\in\mathcal{E}^{1,NA}.$
The usefulness of this inequality is similar to that of the ‘Archimedean’ counterpart $K^{\beta}\leq K$, already discussed. Sending $\beta\to\infty$, and using Theorem 1.2, we obtain that $K\{u_{t}\}\leq K^{NA}(u)$. Since the reverse inequality is already known [61, Theorem 1.2], we obtain an alternative argument for the recently announced YTD theorem of Boucksom–Jonsson:
Theorem 1.4 (=Corollary 9.4).
[23] For any $u\in\mathcal{E}^{1,NA}$, one has
$K\{u_{t}\}=K^{NA}(u).$
Moreover, a unique csck metric exists in $c_{1}(L)$ if and only if there exists $\gamma>0$ such that
$K^{NA}(u)\geq\gamma J^{NA}(u),\ \ u\in\mathcal{H}^{NA}_{\mathcal{M}}.$
This result, together with its more general version allowing holomorphic vector fields obtained in [23], verifies a conjecture of C. Li who proved the existence direction [61].
This YTD theorem and our Theorem 1.1 imply that our notion of uniform $K^{\beta}$-stability is equivalent to uniform $K$-stability for model filtrations, as they are both equivalent to the existence of a unique cscK metric. Of independent interest, at the end of Section 9 we give an alternative direct proof of this equivalence that bypasses cscK metrics.
In light of the preceding discussion and the examples of [2], it is natural to wonder if these stability notions are equivalent to the traditional notion of uniform K-stability as suggested by Székelyhidi [81] and formulated by Dervan–Boucksom–Hisamoto–Jonsson [46, 17]:
Question 1.5.
Is uniform K-stability equivalent to uniform $K^{\beta}$-stability (for $\beta\gg 0$)?
By the inequality $K\geq K^{\beta}$ noted above, $K^{\beta}$-stability implies $K$-stability, mirroring that Ding stability implies $K$-stability in the Fano case [5]. The converse direction remains a subject for further investigation.
To the best of our knowledge, our $K^{\beta}$ functionals have not been previously considered in the K-stability literature. We note, however, that [29, Section 4.1] considered a similar-flavored functional in their alternative proof of the convexity of the $K$ energy. Their construction involves an additional Banach-Saks averaging, together with a different type of quantization in the definition of their $\mathcal{M}_{k}$ functional. At present, it remains unclear whether variants of our $K^{\beta}$ functionals yield other new notions of stability, suggesting another intriguing direction for future investigation.
Although we did not make explicit use of the expansion (5), it provided key motivation for our extensive reliance on transcendental quantization. Such expansions have already played an important role in influential studies of Kähler–Einstein type metrics by Wu and Jeffres-Mazzeo-Rubinstein [89], [59, Section 9] (cf. [77, (6.3)]), where they were employed to initiate a Ricci continuity path from -$\infty$, following a suggestion of Tian–Yau [87].
Our results about the $\beta$-entropy $Ent^{\beta}(\cdot)$ hold for any general compact Kähler manifold, but to stay brief, in this paper we do not study the relationship of this functional with holomorphic vector fields of the manifold. Instead, we refer the reader to [61, 57, 58, 23] for exciting recent developments in this direction.
Acknowledgments.
K.Z. is especially grateful to J. Chu for discussions on the expansion of $u^{\beta}$ back in 2021. We are grateful to P. Mesquita Piccione, J. Han, M. Xia for answering numerous questions, and we thank Y. Liu and Y. Odaka for suggestions that improved the presentation.
T.D. was partially supported by a Sloan Fellowship, Simons Fellowship and NSF grant DMS-2405274. This material is based upon work supported by NSF grant DMS-1928930, while T.D. was in residence at the Simons Laufer Mathematical Sciences Institute during Fall 2024. K.Z. is supported by NSFC grants 12571060, 12271038 and 12271040.
Both authors visited Chalmers University in Spring 2025, where part of this work was completed. We thank the mathematics department for the hospitality.
During the final stage of this project, we became aware of the related work announced by Boucksom–Jonsson [19]. Until then we had not considered model filtrations, but afterwards we naturally wondered about the relationships between their work and ours, resulting in the brief §9, where we give an alternative argument for their Theorem 1.4 using $K^{\beta}$ functionals. We are grateful to them for sharing their preprint [23] and also for related discussions. As far as we can tell, our approaches are fundamentally different. The work of Boucksom–Jonsson is a culmination of the non-Archimedean approach to K-stability, developed over the past decade (see [17, 10, 20, 21] etc.), and relies on an approximation argument at the level of model filtrations/non-Archimedean data, using [4, 48]. In contrast, our approach is ‘Archimedean’. We introduce novel $K^{\beta}$ functionals that approximate the $K$ energy directly via transcendental quantization at the level of potentials (inspired by [6]).
Paralleling our transcendental YTD correspondence Theorem 6.2, we acknowledge an independent work by Mesquita–Piccione and Witt Nyström [66], proving a different kind of transcendental YTD existence theorem, more in the spirit of [61], obtained through the resolution of non-Archimedean Monge–Ampère equations.
2 Preliminaries
We collect some standard results from the literature that will be used in this work. Throughout this paper $(X,\omega)$ denotes a compact Kähler manifold of dimension $n$.
Pluripotential theory.
The space of smooth Kähler potentials is defined as follows:
$\mathcal{H}_{\omega}:=\{u\in C^{\infty}(X)\textup{ s.t. }\omega_{u}:=\omega+\mathrm{dd^{c}}u>0\},$
where $\mathrm{dd^{c}}:=\frac{i}{2\pi}\partial\bar{\partial}.$ The metric completion of $\mathcal{H}_{\omega}$ with respect to the $d_{1}$ metric [32] is denoted by $\mathcal{E}^{1}$, which is called the space of finite energy potentials [56]. More precisely, given $u,v\in\mathcal{E}^{1}$, the $d_{1}$-distance between $u$ and $v$ is given as
$d_{1}(u,v)=I(u)+I(v)-2I(P(u,v)),$
where $I(\cdot)$ denotes the Monge-Ampère energy (recall (3)) and $P(u,v)\in\mathcal{E}^{1}$ is the potential theoretic minimum of $u$ and $v$, i.e., the largest potential in $\mathcal{E}^{1}$ lying below $u$ and $v$. In particular, $d_{1}(0,u)=-I(u)$ when $u\leq 0$.
Given $a,b\in[-\infty,\infty]$ with $a<b$, a path $(a,b)\ni t\mapsto u_{t}\in\mathcal{E}^{1}$ is called a finite energy subgeodesic segment if $U_{z}(x):=u_{-\log|z|^{2}}(x)$ (with $x\in X$ and $z$ being a complex) satisfies
$p_{1}^{*}\omega+\mathrm{dd}^{c}_{X\times{\mathbb{C}}}U_{z}\geq 0$
as a current on $X\times\{e^{-a/2}<|z|<e^{-b/2}\}$, where $p_{1}:X\times{\mathbb{C}}\to X$ is the projection map.
Let $h_{0},h_{1}\in\mathcal{E}^{1}$. As shown in [32], a distinguished $d_{1}$-geodesic $0,1]\ni t\mapsto h_{t}\in\mathcal{E}^{1}_{\omega}$ connecting $h_{0},h_{1}$ can be obtained as Perron envelope of finite energy subgeodesics:
$h_{t}:=\sup\{v_{t}\ |\ [0,1]\ni l\ \mapsto v_{l}\in\mathcal{E}^{1}\textup{ is a subgeodesic with }\limsup_{l\to 0,1}v_{l}\leq h_{0,1}\},\ t\in[0,1].$ (13)
We call (13) the finite energy geodesic connecting $h_{0},h_{1}$.
In case $a=0$ and $b=\infty$ then the curve $(0,\infty)\ni t\mapsto u_{t}\in\mathcal{E}^{1}$ is called a finite energy subgeodesic ray, which we will denote by $\{u_{t}\}$. All rays in this paper will be assumed to emanate from $u_{0}=0\in\mathcal{H}_{\omega}$, meaning that $d_{1}(u_{t},0)\to 0$ as $t\searrow 0$. It is well known that $t\mapsto I(u_{t})$ is convex (resp. affine linear) when $\{u_{t}\}$ is a subgeodesic (resp. geodesic).
A subgeodesic ray $\{u_{t}\}$ is called sublinear if $u_{t}\leq Ct$ for some $C>0$. Recall that a geodesic ray is automatically sublinear due to [40, Theorem 3.7(iv)]. Finally, following the formalism of [37], by $\mathcal{R}^{1}$ we denote the space of finite energy geodesic rays emanating from 0, which is complete with respect to the chordal metric $d_{1}^{c}$ defined in (10) [37, Theorem 1.4].
A geodesic ray $\{u_{t}\}\in\mathcal{R}^{1}$ is bounded if $u_{t}\in\mathcal{E}^{1}\cap L^{\infty},\ t\geq 0$. [33, Theorem 1] implies that such rays are automatically linearly bounded: there exists $C>0$ such that $|u_{t}|\leq Ct,\ t\geq 0$.
For more comprehensive works on the above topics we refer to [34, 37].
Test configurations.
For the rest of this section we assume that $(X,\omega)$ is polarized, with $\omega\in c_{1}(L)$ for some ample line bundle $L$.
[…]
In practice, one often encounters semi-ample test configurations, for which $\mathcal{L}$ is only assumed to be relatively semi-ample. By [17, §2.5], such a generalization is hardly essential. Indeed, semi-ample test configurations are simply pullbacks of their ample models, and both induce the same geodesic ray. As a result, in this work, slightly abusing precision, we think of ample and semi-ample test configurations as the same objects.
[…]
Conversely (see [17, §2.6]), for any semi-ample test configuration $(\mathcal{X},\mathcal{L})$, after replacing $\mathcal{L}$ with $\mathcal{L}+a\mathcal{X}_{0}$ for appropriate $a\in\mathbb{Q}$, the resulting test configuration has the same Futaki invariant, but induces a flag ideal $\mathfrak{a}$ on $X\times{\mathbb{C}}$ such that the sheaf $p_{1}^{*}(mL)\otimes\mathfrak{a}$ is globally generated for some divisible $m>0$. This ideal gives rise to a semi-ample test configuration $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ with a morphism $\mathcal{X}^{\prime}\xrightarrow{\tau}\mathcal{X}$ such that $\tau^{*}\mathcal{L}=\mathcal{L}^{\prime}$.
Most importantly, to any semi-ample test configuration $\mathcal{T}$ one can attach a $C^{1,1}$ geodesic ray $\{u_{t}^{\mathcal{T}}\}$ [72, 73, 76, 5, 30]. When the semi-ample test configuration is obtained via a flag ideal $\mathfrak{a}$ as described above, one further sees that the singularity type of $U^{\mathcal{T}}_{z}:=u^{\mathcal{T}}_{-\log|z|^{2}}$ near $X\times\{0\}$ is comparable to $\mathfrak{a}^{1/m}$, in the sense that
$U^{\mathcal{T}}_{z}=\frac{1}{m}\log\sum_{i}|f_{i}|^{2}+O(1),$ (14)
where $\{f_{i}\}$ is a finite set of local generators of $\mathfrak{a}$. For this reason, $\{u_{t}^{\mathcal{T}}\}$ is also called a geodesic ray with algebraic singularities in the terminology of [10, §4.4].
Demailly regularization.
We recall an important construction going back to [44, 10]. Let $\{u_{t}\}$ be a finite energy subgeodesic ray with $u_{t}\leq 0$. Put $U:=U_{z}=u_{-\log|z|^{2}}$, which extends to a quasi-plurisubharmonic (qpsh) function on $X\times\Delta$. For any $m\geq 1$, let $\mathfrak{a}_{m}:=\mathcal{I}(mU)$ be the multiplier ideal sheaf of $mU$. Note that $\mathfrak{a}_{m}$ is trivial away from $X\times\{0\}$ (by [94]), so it extends to a flag ideal on $X\times{\mathbb{C}}$. As shown in [10, Lemma 5.6], one can find a sufficiently large $m_{0}>0$, depending only on $(X,L)$, such that $p_{1}^{*}((m+m_{0})L)\otimes\mathfrak{a}_{m}$ is globally generated for any $m\geq 1$. Let $\mathcal{X}_{m}\xrightarrow{\mu_{m}}X\times{\mathbb{C}}$ be the normalized blowup along $\mathfrak{a}_{m}$ with exceptional divisor $E_{m}$. Put $\mathcal{L}_{m}:=\mu_{m}^{*}p_{1}^{*}L-\frac{1}{m+m_{0}}E_{m}$. Then $\mathcal{T}_{m}:=(\mathcal{X}_{m},\mathcal{L}_{m})$ is a semi-ample test configurations, inducing a geodesic ray $\{u_{t}^{\mathcal{T}_{m}}\}$. Using (14), one then observes that
$\int_{X\times\Delta}e^{(m+m_{0})U^{\mathcal{T}_{m}}-mU}\omega^{n}\wedge idz\wedge d\bar{z}<\infty.$ (15)
Based on this simple fact, a remarkable result was obtained by Li [61]: if $\{u_{t}\}\in\mathcal{R}^{1}$ is a geodesic ray with $Ent\{u_{t}\}<\infty$, then it is maximal in the sense of [10, Definition 6.5]. More specifically, we have that (cf. [10, Lemma 5.7], [75, Theorem 4.4.1])
$\{u_{t}^{\mathcal{T}_{m}}\}\xrightarrow{d_{1}^{c}}\{u_{t}\}\text{ as }m\to\infty.$
3 Transcendental quantization
Let $(X,\omega)$ be a compact Kähler manifold of dimension $n$. Recall that for any $\beta>0$ and $u\in\mathcal{E}^{1}$, $e^{-\beta u}\in L^{p}$ for any $p>1$ [94]. Thus, by the theory of Monge-Ampère equations [9, Theorem C], there exists a unique continuous $\omega$-psh function $u^{\beta}$ satisfying
$(\omega+\mathrm{dd^{c}}u^{\beta})^{n}=e^{\beta(u^{\beta}-u)}\omega^{n}.$ (16)
Based on the speculations of [6, Section 3.1], we call $u^{\beta}$ the transcendental quantization of $u$ at level $\beta$ (see [34, Appendix A.3] for a short survey on this topic).
There are some immediate properties of $u^{\beta}$, reminiscent of Kähler quantization:
Lemma 3.1.
One has
(i) $(u+C)^{\beta}=u^{\beta}+C$ for any $u\in\mathcal{E}^{1}$ and $C\in\mathbb{R}$.
(ii) $u\geq v$ implies that $u^{\beta}\geq v^{\beta}$.
(iii) $Ent(u^{\beta})<\infty$ for any $u\in\mathcal{E}^{1}$.
Proof.
The first point is trivial and the second follows from the comparison principle [6, Lemma 2.5] (whose proof applies to potentials of $\mathcal{E}^{1}$). Finally, we argue the last point:
$\begin{split}Ent(u^{\beta})&=\frac{1}{V}\int_{X}\log\frac{\omega^{n}_{u^{\beta}}}{\omega^{n}}\omega^{n}_{u^{\beta}}=\frac{\beta}{V}\int_{X}(u^{\beta}-u)\omega^{n}_{u^{\beta}}\leq\frac{\beta}{V}\int_{X}|u^{\beta}-u|\omega^{n}_{u^{\beta}}\leq\beta C_{n}d_{1}(u^{\beta},u)<\infty.\end{split}$
The preceding observation indicates that $u^{\beta}$ serves as a regularization of $u$, particularly in cases where $u\in\mathcal{E}^{1}_{\omega}$ and $\operatorname{Ent}(u)=\infty$. We now proceed to show that $u^{\beta}\xrightarrow{d_{1}}u$.
Lemma 3.2.
For any $u\in\mathcal{E}^{1}$, one has $u^{\beta}\xrightarrow{d_{1}}u$ as $\beta\to\infty$.
Proof.
We first show that $u^{\beta}\xrightarrow{a.e.}u$ as $\beta\to\infty$. Let $u\in\mathcal{E}^{1}$ and $\delta>1$. Consider $u_{\delta}:=\delta^{-1}P(\delta u)\in\mathcal{E}^{1}$ as in [38, Proposition 2.15] (Proposition 2.18 in the published version). Then $u_{\delta}\leq u\text{ and }\omega+\mathrm{dd^{c}}u_{\delta}\geq(\frac{\delta-1}{\delta})\omega.$ Thus:
$(\omega+\mathrm{dd^{c}}u_{\delta})^{n}\geq e^{\beta((u_{\delta}+\frac{n}{\beta}\log(\frac{\delta-1}{\delta}))-u_{\delta})}\omega^{n}\ \ \ \textup{ and }\ \ \ (\omega+\mathrm{dd^{c}}u^{\beta})^{n}\leq e^{\beta(u^{\beta}-u_{\delta})}\omega^{n},$
Applying [6, Lemma 2.5] again, we obtain that
$u^{\beta}\geq u_{\delta}+\frac{n}{\beta}\log\Big(\frac{\delta-1}{\delta}\Big).$ (17)
In particular, $\liminf_{\beta\to\infty}u^{\beta}(x)\geq u_{\delta}(x),\ \text{for all }\delta>1,x\in X.$ Since $u_{\delta}\xrightarrow{a.e.}u$ as $\delta\searrow 1$ ([38, Proposition 2.15]), we obtain that
$\liminf_{\beta\to\infty}u^{\beta}(x)\geq u(x)\text{ for almost all }x\in X.$
We now bound $u^{\beta}$ from above. Pick a smooth sequence $u_{k}\searrow u$ pointwise. Using the fact that $u_{k}^{\beta}\geq u^{\beta}$ and that $u_{k}^{\beta}\xrightarrow{C^{0}}u_{k}$ as $\beta\to\infty$ ([31, Theorem 1.1], [34, Theorem A.10]),
$\limsup_{\beta\to\infty}u^{\beta}(x)\leq\lim_{\beta\to\infty}u_{k}^{\beta}(x)=u_{k}(x)\text{ for any }k\text{ and }x\in X.$
Thus, $\limsup_{\beta\to\infty}u^{\beta}(x)\leq u(x),x\in X$. So $u^{\beta}\xrightarrow{a.e.}u$ as $\beta\to\infty$, as claimed.
Let $\delta>1$. We estimate $d_{1}(u^{\beta},u)$ as follows:
$\begin{split}d_{1}(u,u^{\beta})&\leq d_{1}(u_{\delta}+\frac{n}{\beta}\log\Big(\frac{\delta-1}{\delta}\Big),u^{\beta})+d_{1}(u,u_{\delta})-\frac{n}{\beta}\log\Big(\frac{\delta-1}{\delta}\Big)\\ &=I(u^{\beta})-I(u_{\delta}+\frac{n}{\beta}\log\Big(\frac{\delta-1}{\delta}\Big))+d_{1}(u,u_{\delta})-\frac{n}{\beta}\log\Big(\frac{\delta-1}{\delta}\Big)\\ &\leq\frac{1}{V}\int_{X}|u^{\beta}-u_{\delta}|\omega^{n}_{u_{\delta}}+d_{1}(u,u_{\delta})-\frac{2n}{\beta}\log\Big(\frac{\delta-1}{\delta}\Big).\end{split}$
Sending $\beta\to\infty$, using dominated convergence and [32, Theorem 3],
$\limsup_{\beta\to\infty}d_{1}(u,u^{\beta})\leq\frac{1}{V}\int_{X}|u-u_{\delta}|\omega^{n}_{u_{\delta}}+d_{1}(u,u_{\delta})\leq C_{n}d_{1}(u,u_{\delta}).$
Finally, sending $\delta\searrow 1$ and using [38, Proposition 2.15], we conclude. ∎
Next we show convergence of the entropy functional under transcendental quantization:
Lemma 3.3.
For any $u\in\mathcal{E}^{1}$, $Ent(u^{\beta})\to Ent(u)$ as $\beta\to\infty$.
Proof.
If $Ent(u)=\infty$, then trivially $Ent(u)\geq Ent(u^{\beta})$. If $Ent(u)<\infty$, we can use Jensen’s inequality to derive that
[…]
Thus $Ent(u)\geq Ent(u^{\beta})$ for any $u\in\mathcal{E}^{1}$. So we conclude since $Ent(\cdot)$ is $d_{1}$-lsc ([8], cf. [34, Corollary 4.40]). ∎
The following simple but important estimate can be extracted from the above proof.
Corollary 3.4.
For any $u\in\mathcal{E}^{1}$,
$Ent(u)\geq\beta(I(u^{\beta})-I(u))\geq Ent(u^{\beta}).$
The next result says that our quantization procedure improves convergence:
Lemma 3.5.
Assume that $u_{i}\xrightarrow{L^{1}}u$ in $\mathcal{E}^{1}$, then $u^{\beta}_{i}-u^{\beta}||_{C^{0}}\to 0$ for any $\beta>0$. In particular, $u_{i}^{\beta}\xrightarrow{d_{1}}u^{\beta}$.
Proof.
Since for any $p>0$, $e^{-pu_{i}}$ and $e^{-pu}$ are integrable by [94], [45, Theorem 0.2] implies that $\int_{X}|e^{-pu_{i}}-e^{-pu}|\omega^{n}\to 0.$ In particular, there exists $A_{p}>0$ such that
[…]
On the other hand, by the stability estimate [55, Remark 3.2], for any $\beta>0$,
[…]
where $C$ depends on $X,\omega,p,A_{p}$ for some $p>\beta$. We then conclude since $e^{-\beta u_{i}}\xrightarrow{L^{1}}e^{-\beta u}.$
The $\beta$-entropy and quantization.
For $u\in\mathcal{E}^{1}$ and $\beta>0$, the $\beta$-entropy $Ent^{\beta}(u)$ is defined by extending (4):
$Ent^{\beta}(u):=\sup_{v\in\mathcal{E}^{1}}\left(-\log\int_{X}e^{\beta(v-u)}\frac{\omega^{n}}{V}+\beta(I(v)-I(u))\right).$ (18)
We link this functional to our quantization scheme:
Proposition 3.6.
For any $u\in\mathcal{E}^{1}$ and $\beta>0$,
$Ent^{\beta}(u)=\beta(I(u^{\beta})-I(u)).$
Proof.
That $Ent^{\beta}(u)\geq\beta(I(u^{\beta})-I(u))$ follows from (18) and (16). Conversely, for $v\in\mathcal{E}^{1}_{\omega}$ we have
[…]
This completes the proof. ∎
The above argument shows that $u^{\beta}$ attains the supremum in (18). Moreover, taking $\mathcal{H}_{\omega}\ni u_{i}\searrow u$ and applying Lemma 3.5, then $u_{i}^{\beta}\in\mathcal{H}_{\omega}$ is a smooth maximizing sequence for (18), showing that (4) holds for $u\in\mathcal{E}^{1}$ as well.
We also observe that $\beta\mapsto Ent^{\beta}(\cdot)$ satisfies certain concavity.
Corollary 3.7.
For any $0<\beta<\beta^{\prime}$ and $u\in\mathcal{E}^{1}$, one has $\frac{Ent^{\beta}(u)}{\beta}\geq\frac{Ent^{\beta^{\prime}}(u)}{\beta^{\prime}}.$
Proof.
Using (18), we can write
$\begin{split}Ent^{\beta}(u)&\geq-\log\int_{X}e^{\beta(u^{\beta^{\prime}}-u)}\frac{\omega^{n}}{V}+\beta(I(u^{\beta^{\prime}})-I(u))\geq 0+\frac{\beta}{\beta^{\prime}}Ent^{\beta^{\prime}}(u),\end{split}$
where in the last line we used Hölder’s inequality and that $\int_{X}e^{\beta^{\prime}(u^{\beta^{\prime}}-u)}\omega^{n}/V=1$. This completes the proof. ∎
As a direct consequence of the identity $Ent^{\beta}(u)=\beta(I(u^{\beta})-I(u))$, we now derive several properties of the associated $\beta$-entropy.
Lemma 3.8.
One has
(i) $Ent(u)\geq Ent^{\beta}(u)\geq Ent(u^{\beta})\geq 0$ for any $u\in\mathcal{E}^{1}$ and $\beta>0$.
(ii) $Ent^{\beta}(u)\to Ent(u)$ as $\beta\to\infty$ for any $u\in\mathcal{E}^{1}$.
(iii) $\frac{1}{\beta}Ent^{\beta}(u)\to 0$ as $\beta\to\infty$ for any $u\in\mathcal{E}^{1}$.
Proof.
These follow from Lemma 3.3, Corollary 3.4 and Lemma 3.5. ∎
Compared to the usual entropy, our $Ent^{\beta}(\cdot)$ has better continuity properties on $\mathcal{E}^{1}$:
Lemma 3.9.
Let $\beta>0.$ The following holds.
(i) $Ent^{\beta}(\cdot)$ is $L^{1}$-lsc. Namely, if $u_{i}\xrightarrow{L^{1}}u$ in $\mathcal{E}^{1}$ then $Ent^{\beta}(u)\leq\liminf_{i\to\infty}Ent^{\beta}(u_{i}).$
(ii) $Ent^{\beta}(\cdot)$ is $d_{1}$-continuous. Namely, if $u_{i}\xrightarrow{d_{1}}u$ then $Ent^{\beta}(u)=\lim_{i\to\infty}Ent^{\beta}(u_{i}).$
Proof.
The first point follow from Lemma 3.5 and that $I(\cdot)$ is $L^{1}$-usc [8] (see also [34, Corollary 4.14]). Then second follows from Lemma 3.5 as well. ∎
Another interesting feature of the quantized entropy $Ent^{\beta}(\cdot)$ is that it is monotone with respect to the parameter $\beta$.
Proposition 3.10.
Let $0<\beta_{1}\leq\beta_{1}$. Then for any $u\in\mathcal{E}^{1}$, $Ent^{\beta_{1}}(u)\leq Ent^{\beta_{2}}(u).$
Proof.
By approximation, it suffices to prove the monotonicity for $u\in\mathcal{H}_{\omega}$. In this case, $\{u^{\beta}\}_{\beta\in(0,\infty)}$ forms a smooth family of Kähler potentials. Therefore it suffices to show that
$\frac{d}{d\beta}Ent^{\beta}(u)=I(u^{\beta})-I(u)+\frac{\beta}{V}\int_{X}\frac{du^{\beta}}{d\beta}\omega^{n}_{u^{\beta}}\geq 0.$
To this end, we differentiate (16) with respect to $\beta$:
$\Delta_{\omega_{u^{\beta}}}\frac{du^{\beta}}{d\beta}=(u^{\beta}-u)+\beta\frac{du^{\beta}}{d\beta}.$
Integrating against $\omega^{n}_{u^{\beta}}$, we derive that $\frac{\beta}{V}\int_{X}\frac{du^{\beta}}{d\beta}\omega^{n}_{u^{\beta}}=-\frac{1}{V}\int_{X}(u^{\beta}-u)\omega^{n}_{u^{\beta}}.$ Thus,
$I(u^{\beta})-I(u)+\frac{\beta}{V}\int_{X}\frac{du^{\beta}}{d\beta}\omega^{n}_{u^{\beta}}=I(u^{\beta})-I(u)-\frac{1}{V}\int_{X}(u^{\beta}-u)\omega^{n}_{u^{\beta}}\geq 0,$
finishing the proof. ∎
As a consequence of the above result, Lemma 3.3 and Corollary 3.4 we note the following:
Corollary 3.11.
For $u\in\mathcal{E}^{1}$ we have $Ent(u^{\beta})\leq Ent^{\beta}(u)\leq Ent(u)$ and $Ent^{\beta}(u)\nearrow Ent(u)$ as $\beta\nearrow\infty$.
We also record a simple estimate, which will be used later.
Lemma 3.12.
Let $u\in\mathcal{E}^{1}$ with $u\leq 0$. Then $d_{1}(0,u^{\beta})\leq d_{1}(0,u)$ for any $\beta>0$.
Proof.
By Lemma 3.1(ii) we have $u^{\beta}\leq 0$. Thus by Proposition 3.6,
[…]
where we used $Ent^{\beta}(u)\geq 0$ (see Lemma 3.8(i)). ∎
Păun’s positivity result.
In the classical setting of algebraic quantization, a useful positivity result due to Berndtsson roughly says that the quantization of a subgeodesic is still a subgeodesic [13, Proposition 3.1], a crucial point in the second author’s alternative proof of the YTD conjecture for Kähler-Einstein type metrics [96]. Later in §7, a transcendental replacement for Berndtsson’s result will be needed. It is given by the following positivity result that can be extracted from Păun’s paper [74, §3] (cf. also [29, Theorem 4.1]).
Theorem 3.13.
Let $\Sigma:=\{a<|z|<b\}\subset{\mathbb{C}}$ be an annulus. Let $X$ be a compact Kähler manifold of dimension $n$ and $\xi$ be a Kähler class on $X$. Let $\Omega$ be a smooth positive volume form on $X$. We denote by $\operatorname{Ric}(\Omega):=-\mathrm{dd^{c}}\log\det\Omega$ the Ricci form associated with $\Omega$. Denote by $\pi$ the projection map $X\times\Sigma\to X$. Assume that there exists a smooth semipositive form $\mathcal{A}$ on $X\times\Sigma$ in the $(1,1)$-class $\pi^{*}\xi$ such that $\mathcal{A}|_{X\times\{z\}}-Ric(\Omega)$ is a Kähler form on $X\times\{z\}$ for all $z\in\Sigma$. For any $z\in\Sigma$, consider the unique smooth solution $\Phi_{z}$ of the Aubin–Yau type equation on $X\times\{z\}$:
$(\mathcal{A}|_{X\times\{z\}}-\operatorname{Ric}(\Omega)+\mathrm{dd}_{X}^{c}\Phi_{z})^{n}=e^{\Phi_{z}}\Omega.$
Then $\mathcal{A}-\pi^{*}\operatorname{Ric}(\Omega)+\mathrm{dd}^{c}_{X\times\Sigma}\Phi_{z}$ is a smooth semipositive $(1,1)$-form on $X\times\Sigma.$
Actually in [74] Păun proved a much more general positivity result for a family of Kähler manifolds. But for us the above simple product case is just enough.
Let us explain how to apply the above result to our setting. First assume that $\{u_{t}\}_{t\in(0,1)}$ is a smooth subgeodesic segment in $\mathcal{E}^{1}$, i.e., under the coordinate change $t:=-\log|z|^{2}$,
$\pi^{*}\omega+\mathrm{dd}^{c}_{X\times\Sigma}u_{-\log|z|^{2}}\geq 0\ \textup{ on }\ X\times\Sigma,$
where $\Sigma:=\{e^{-1}<|z|^{2}<1\}\subset{\mathbb{C}}$. Assume further that there exists $\eta>0$ such that
$\pi^{*}\omega+\mathrm{dd^{c}}_{X\times\Sigma}{u_{-\log|z|^{2}}}\geq\eta\pi^{*}\omega.$
Then for large enough $\beta>0$, $\beta\omega_{u_{t}}+\operatorname{Ric}(\omega)>0.$ So putting
$\mathcal{A}:=\pi^{*}\beta\omega+\mathrm{dd}^{c}_{X\times\Sigma}\beta u_{-\log|z|^{2}}+\pi^{*}\operatorname{Ric}(\omega),$
we obtain a smooth semipositive $(1,1)$-form on $X\times\Sigma$ whose restriction on each fiber $X\times\{z\}$ is the Kähler form $\beta\omega_{u_{t}}+\operatorname{Ric}(\omega)$. Let
$\Phi_{z}:=\beta u_{-\log|z|^{2}}^{\beta}-\beta u_{-\log|z|^{2}}.$
Then the Monge–Ampère equation of $u_{t}^{\beta}$ (16) translates to
$(\mathcal{A}|_{X\times\{z\}}-\operatorname{Ric}(\omega)+\mathrm{dd}_{X}^{c}\Phi_{z})^{n}=e^{\Phi_{z}}\beta^{n}\omega^{n}$
for any $z\in\Sigma$. Denoting $\Omega:=\beta^{n}\omega^{n}$, we have $\operatorname{Ric}(\Omega)=\operatorname{Ric}(\omega)$. So Theorem 3.13 applies, giving us that
$\pi^{*}\omega+\mathrm{dd}^{c}_{X\times\Sigma}u^{\beta}_{-\log|z|^{2}}\geq 0,$
i.e., $\{u_{t}^{\beta}\}_{t\in(0,1)}$ is a subgeodesic segment.
By approximation, the above consideration extends to a general subgeodesic segment in $\mathcal{E}^{1}$, as shown in the next result.
Proposition 3.14.
Assume that $(0,1)\ni t\mapsto u_{t}\in\mathcal{E}^{1}$ is subgeodesic segment. Suppose that $\operatorname{Ric}\omega\geq-A\omega$ for some $A\geq 0$. Then for any $\eta\in(2A/\beta,1)$ and $\beta>2A$ the curve $(0,1)\ni t\mapsto((1-\eta)u_{t})^{\beta}\in\mathcal{E}^{1}_{\omega}$ is subgeodesic segment as well.
Proof.
Assume without loss of generality that $u_{t}\leq 0$. For any $\varepsilon\in(0,\eta)$ put
$u_{t,\varepsilon}:=(1-\eta)u_{t}+\frac{\varepsilon}{e^{t}}.$
Consider the projection $X\times\{e^{-1}<|z|^{2}<1\}\xrightarrow{\pi}X$ and $t=-\log|z|^{2}$. Then
$\pi^{*}\omega+\mathrm{dd^{c}}u_{-\log|z|^{2},\varepsilon}\geq\eta\pi^{*}\omega+\varepsilon idz\wedge d\bar{z}>0.$
Thus $\{u_{t,\varepsilon}\}$ is subgeodesic and also it is clear that $u_{t,\varepsilon}\searrow(1-\eta)u_{t}$ pointwise as $\varepsilon\searrow 0$. Therefore, it suffices to show that $\{u_{t,\varepsilon}^{\beta}\}$ is subgeodesic (note from Lemma 3.5 that $u_{i}\xrightarrow{L^{1}}u$ implies $u_{i}^{\beta}\xrightarrow{C^{0}}u^{\beta}$).
To this end, we need to further regularize $u_{t,\varepsilon}$. For any $0<\delta<\varepsilon$, by [15, Theorem 2] we can find smooth approximating functions $U_{\varepsilon,\delta}$ on $X\times\{e^{-1}+\delta<|z|^{2}<1-\delta\}$ such that
$\pi^{*}\omega+\mathrm{dd^{c}}U_{\varepsilon,\delta}\geq\frac{2A}{\beta}\pi^{*}\omega+(\varepsilon-\delta)idz\wedge d\bar{z}>0.$
Moreover (by averaging), we can arrange that $U_{\varepsilon,\delta}$ is $S^{1}$-invariant in $z$, yielding a smooth subgeodesic $\{u_{t,\varepsilon,\delta}\}\subset\mathcal{H}_{\omega}$ with $u_{t,\varepsilon,\delta}\searrow u_{t,\varepsilon}$ pointwise as $\delta\searrow 0$.
It follows that
$\beta\pi^{*}\omega+\mathrm{dd^{c}}\beta U_{\varepsilon,\delta}+\operatorname{Ric}\omega\geq A\pi^{*}\omega+\beta(\varepsilon-\delta)idz\wedge d\bar{z}>0.$
Therefore, it suffices that argue that $\{u^{\beta}_{t,\varepsilon,\delta}\}$ is subgeodesic for all $\beta>2A$, which now follows Theorem 3.13, and the discussion following it. ∎
Remark 3.15.
If $\operatorname{Ric}(\omega)\geq 0$, then we can take $\eta=0$ (using a limit process $\eta\searrow 0$) in the above result. So in this case $t\mapsto u_{t}$ is subgeodesic for any $\beta>0$ whenever $t\mapsto u_{t}$ is.
4 Properness of the $K^{\beta}$ energy
In this section we turn to the analysis of the K energy and $K^{\beta}$ energy. Recall from (2) that
$K(u):=Ent(u)-\mathcal{J}_{\operatorname{Ric}(\omega)}(u),\ u\in\mathcal{E}^{1}\ ,$
where for some smooth closed $(1,1)$-form $\chi$ we define
$\mathcal{J}_{\chi}(u):=nI_{\chi}(u)-\bar{\chi}I(u)=\frac{1}{V}\int_{X}u\chi\wedge\sum_{i=0}^{n-1}\omega^{i}\wedge\omega^{n-1-i}_{u}-\bar{\chi}I(u),\ \bar{\chi}=\frac{n}{V}\int_{X}\chi\wedge\omega^{n-1}.$
For $\beta>0$, recall from (7) the definition of our $K^{\beta}$ energy:
$K^{\beta}(u):=Ent^{\beta}(u)-\mathcal{J}_{\operatorname{Ric}(\omega)}(u).$
As a consequence of Proposition 3.10 we note the following fact for $u\in\mathcal{E}^{1}$:
$K^{\beta}(u)\nearrow K(u)\ \textup{ as }\ \beta\nearrow\infty$ (19)
In our first technical result we show that the distance $d_{1}(u,u^{\beta})$ is comparable to the growth of $\frac{1}{\beta}Ent^{\beta}(u)$. The role of this result in our work is analogous with the role of the partial $C^{0}$ estimate from the Fano case, first proposed by Tian (cf. [84, (0.2)]).
Proposition 4.1.
There exists $C_{n}>0$ such that for any $u\in\mathcal{E}^{1}$ with $u\leq 0$ and $\beta>1,$
$\frac{1}{\beta}Ent^{\beta}(u)\leq d_{1}(u^{\beta},u)\leq\frac{1}{\beta}\left(Ent^{\beta}(u)+C_{n}\log(d_{1}(0,u)+2)+C_{n}\log\beta\right).$
Proof.
For any $\delta\in(1,2)$, let $u_{\delta}:=\delta^{-1}P(\delta u)$. Then, using (17) and $u\leq 0$
$u^{\beta}\geq u_{\delta}+\frac{n}{\beta}\log\Big(\frac{\delta-1}{\delta}\Big)\geq P(\delta u)+\frac{n}{\beta}\log\Big(\frac{\delta-1}{\delta}\Big).$
Thus,
[…]
In the last line we used $I(u)-I(v)\leq d_{1}(u,v)$.
We control $d_{1}(P(\delta u),u)$ as follows (using [32, Theorem 3] and [35, Lemma 4.4]):
$\begin{split}d_{1}(P(\delta u),u)&\leq\frac{C_{n}}{V}\int_{X}|u-P(\delta u)|\omega^{n}_{P(\delta u)}\leq\frac{C_{n}\delta^{n}(\delta-1)}{V}\int_{X}|u|\omega^{n}_{u}\leq C_{n}^{\prime}\delta^{n}(\delta-1)d_{1}(0,u).\end{split}$ (20)
So we arrive at
$d_{1}(u^{\beta},u)\leq\frac{1}{\beta}Ent^{\beta}(u)+2C_{n}^{\prime}\delta^{n}(\delta-1)d_{1}(0,u)-\frac{2n}{\beta}\log\frac{\delta-1}{\delta}.$
Next, we choose $\delta$ such that $(\delta-1)(d_{1}(0,u)+1)=\frac{1}{\beta}.$ Since $\beta>1$, such $\delta$ satisfies $1<\delta<2$, and we have
$d_{1}(u^{\beta},u)\leq\frac{1}{\beta}Ent^{\beta}(u)+\frac{2^{n+1}C_{n}^{\prime}}{\beta}+\frac{2n}{\beta}\log(2\beta d_{1}(0,u)+2\beta).$
This establishes the upper bound for $d_{1}(u^{\beta},u)$.
Finally, we also have $d_{1}(u^{\beta},u)\geq I(u^{\beta})-I(u)=\frac{1}{\beta}Ent^{\beta}(u)$, finishing the proof. ∎
Proposition 4.2.
Let $\chi$ be a closed smooth $(1,1)$-form. There exists a constant $C>0$, depending only on $X,\omega,\chi,n$ such that
$|\mathcal{J}_{\chi}(u^{\beta})-\mathcal{J}_{\chi}(u)|\leq\frac{C}{\beta^{1/2^{n}}}(Ent^{\beta}(u)+d_{1}(0,u)+1+\log\beta)$
for any $\beta>1$ and $u\in\mathcal{E}^{1}$ with $u\leq 0$.
Proof.
In what follows, the constant $C>0$ only depends on $X,\omega,\chi,n$, but may change from line to line.
Pick $\delta\in(1,2)$. Then, as before, $u\geq P(\delta u)+\frac{n}{\beta}\log\frac{\delta-1}{\delta}$ and $u^{\beta}\geq P(\delta u)+\frac{n}{\beta}\log\frac{\delta-1}{\delta}.$
Let $C>0$ such that $C\omega\leq\chi\leq C\omega$. We then have
[…]
We bound the terms on the last line one by one. By [10, Lemma A.2], we have
$\int_{X}(u^{\beta}-u)\omega\wedge\omega^{n-i-1}_{u}\wedge\omega_{u^{\beta}}^{i}\leq Cd_{1}(u^{\beta},u)^{1/2^{n}}M^{1-1/2^{n}},$
where $M:=\max\{\mathcal{I}(u),\mathcal{I}(u^{\beta})\},$ and $\mathcal{I}(u):=\frac{1}{V}\int_{X}u(\omega^{n}-\omega^{n}_{u})$ denotes Aubin’s $\mathcal{I}$-functional. We assumed that $u\leq 0$, so $u^{\beta}\leq 0$ as well. Due to [32, Theorem 3],
$M\leq(n+1)\max\{J(u),J(u^{\beta})\}\leq(n+1)\max\{d_{1}(0,u),d_{1}(0,u^{\beta})\}=(n+1)d_{1}(0,u),$
where we used Lemma 3.12(i) in the last equality. Therefore,
$\int_{X}(u^{\beta}-u)\omega\wedge\omega^{n-i-1}_{u}\wedge\omega_{u^{\beta}}^{i}\leq Cd_{1}(u^{\beta},u)^{1/2^{n}}d_{1}(0,u)^{1-1/2^{n}}.$ (21)
Similarly, we have
$\int_{X}(u-P(\delta u))\omega\wedge\omega^{n-i-1}_{u}\wedge\omega_{u^{\beta}}^{i}\leq Cd_{1}(u,P(\delta u))^{1/2^{n}}(\max\{d_{1}(0,u),d_{1}(0,P(\delta u))\})^{1-1/2^{n}}.$
Using $d_{1}(u,P(\delta u))\leq C(\delta-1)d_{1}(0,u)$ (recall (20)), we obtain
$d_{1}(0,P(\delta u))\leq d_{1}(0,u)+d_{1}(u,P(\delta u))\leq Cd_{1}(0,u),$
which further implies that
$\int_{X}(u-P(\delta u))\omega\wedge\omega^{n-i-1}_{u}\wedge\omega_{u^{\beta}}^{i}\leq C(\delta-1)^{1/2^{n}}d_{1}(0,u).$
Putting these estimates together, we arrive at
[…]
Now we choose $\delta\in(1,2)$ such that $(\delta-1)^{1/2^{n}}=\frac{1}{\beta},$ which implies that
[…]
To control the terms involving $d_{1}(u^{\beta},u)$, we apply Proposition 4.1 to get
$d_{1}(u^{\beta},u)^{1/2^{n}}d_{1}(0,u)^{1-1/2^{n}}$ $\leq\frac{C}{\beta^{1/2^{n}}}\left(Ent^{\beta}(u)+\log(d_{1}(0,u)+2)+\log\beta\right)^{1/2^{n}}d_{1}(0,u)^{1-1/2^{n}}$
$\leq\frac{C}{\beta^{1/2^{n}}}\left(Ent^{\beta}(u)+d_{1}(0,u)+1+\log\beta\right)^{1/2^{n}}d_{1}(0,u)^{1-1/2^{n}}$ (22)
$\leq\frac{C}{\beta^{1/2^{n}}}\left(Ent^{\beta}(u)+d_{1}(0,u)+1+\log\beta\right).$
Lastly, since $\beta>1$, Proposition 4.1 again implies that
$\begin{split}d_{1}(u^{\beta},u)&\leq\frac{C}{\beta}(Ent^{\beta}(u)+\log(d_{1}(0,u)+2)+\log\beta)\leq\frac{C}{\beta^{1/2^{n}}}(Ent^{\beta}(u)+d_{1}(0,u)+1+\log\beta).\end{split}$
Putting everything together, our result follows. ∎
Lemma 4.3.
There exists a constant $C>0$, depending only on $X,\omega,n$ such that for any $u\in\mathcal{E}^{1}$ with $u\leq 0$,
$\left|\sup_{X}u^{\beta}-\sup_{X}u\right|\leq C+\frac{C}{\beta^{1/2^{n}}}(Ent^{\beta}(u)+d_{1}(0,u)+1+\log\beta).$
Proof.
First, it is well known that (see e.g. [34, Lemma 3.45]) $|\sup_{X}u-\frac{1}{V}\int_{X}u\omega^{n}|\leq C.$ Therefore, it suffices to argue that
$\Big|\int_{X}(u^{\beta}-u)\omega^{n}\Big|\leq\frac{C}{\beta^{1/2^{n}}}(Ent^{\beta}(u)+d_{1}(0,u)+1+\log\beta).$ (23)
Using [10, Lemma A.2] exactly as in the proof of the previous argument we obtain that
$\int_{X}(u^{\beta}-u)\omega^{n}\leq Cd_{1}(u^{\beta},u)^{1/2^{n}}d_{1}(0,u)^{1-1/2^{n}}.$
By (4) we conclude (23). ∎
Recall that an energy functional $G$ defined on $\mathcal{E}^{1}$ is proper (or coercive) if there exists $\gamma,\gamma^{\prime}>0$ such that
$G(u)\geq\gamma J(u)-\gamma^{\prime},\ \ u\in\mathcal{E}^{1}.$
Since $J(u+c)=J(u),\ c\in\mathbb{R}$, this is equivalent to
$G(u)\geq\gamma d_{1}(0,u)-\gamma^{\prime},\ \ u\in\mathcal{E}^{1}\textup{ with }\sup_{X}u=0.$
Indeed if $\sup_{X}u=0$ then
$J(u)\leq d_{1}(0,u)\leq J(u)+C,$ (24)
because $d_{1}(0,u)=-I(u)=J(u)-\int_{X}u\omega^{n},$ and it is well known that $\sup_{X}u-\int_{X}u\omega^{n}$ is always uniformly bounded (cf. [39, Proposition 5.5]).
We now prove the main result of this section, pointing out a strong relationship between properness of the $K$ energy and $K^{\beta}$ energies.
Theorem 4.4.
The following are equivalent:
(i) $K^{\beta}$ energy is proper for some $\beta>1$.
(ii) $K$ energy is proper.
(iii) $X$ admits a unique cscK metric in $\{\omega\}$.
Proof.
Due to [26, 27], it suffices to show that (ii) implies (i). Assume that for some $\delta,C_{0}>0$,
[…]
We show that $K^{\beta}$ is proper as well for all $\beta\gg 1$.
Let $u\in\mathcal{E}^{1}$ with $\sup_{X}u=0$. Using Lemma 3.8, we start a chain of inequalities:
[…]
where we have repeatedly used the estimates of Proposition 4.1, Proposition 4.2 and Lemma 4.3. Reorganizing terms in the above estimate, we arrive at
$\begin{split}&\left(1+\frac{(1+2\delta)C}{\beta^{1/2^{n}}}\right)K^{\beta}(u)+\frac{(1+2\delta)C}{\beta^{1/2^{n}}}\mathcal{J}_{\operatorname{Ric}(\omega)}(u)\\ &\geq\left(\delta-\frac{(1+2\delta)C}{\beta^{1/2^{n}}}\right)d_{1}(0,u)-C-\frac{(1+2\delta)C}{\beta^{1/2^{n}}}\log\beta.\end{split}$
Finally, using that $|\mathcal{J}_{\operatorname{Ric}(\omega)}(u)|\leq Cd_{1}(0,u)$ ([36, Proposition 2.5]) and choosing $\beta$ to be sufficiently large (depending only on $X,\omega,n,\delta$), we obtain $\delta^{\prime}>0$ and $C_{0}^{\prime}>0$ such that $K^{\beta}(u)\geq\delta^{\prime}d_{1}(0,u)-C_{0}^{\prime}$ for any $u\in\mathcal{E}^{1}$ with $\sup_{X}u=0$. This completes the proof. ∎
The following estimate can be extracted from the above proof:
Corollary 4.5.
For any $\beta>1$, there exists a constant $\varepsilon_{\beta}>0$, depending only on $X,\omega,n,\beta$ such that $\varepsilon_{\beta}\to 0$ as $\beta\to\infty$, and that
$(1+\varepsilon_{\beta})K^{\beta}(u)\geq K(u^{\beta})-\varepsilon_{\beta}d_{1}(0,u)-\varepsilon_{\beta},\ \ u\in\mathcal{E}^{1}\ \textup{ with }\ \sup_{X}u=0.$
5 Continuity of the radial $\beta$-entropy
Given $\{u_{t}\}\in\mathcal{R}^{1}$, we define the radial $\beta$-entropy, as follows:
$Ent^{\beta}\{u_{t}\}:=\liminf_{t\to\infty}\frac{Ent^{\beta}(u_{t})}{t}.$ (25)
By Lemma 3.12, one has $Ent^{\beta}(u)\leq\beta d_{1}(u,u^{\beta})\leq 2\beta d_{1}(0,u)$ for any $u\in\mathcal{E}^{1}$ with $\sup_{X}u=0$, from which it follows easily that
$Ent^{\beta}\{u_{t}\}<\infty\text{ for any }\{u_{t}\}\in\mathcal{R}^{1}.$ (26)
To start, we prove the following $d_{1}$-stability result for our transcendental quantization.
Proposition 5.1.
Let $u_{0},u_{1}\in\mathcal{E}^{1}$, $u_{0},u_{1}\leq 0$. For some $C_{n}>0$ the following holds:
$d_{1}(u_{0}^{\beta},u^{\beta}_{1})\leq C_{n}d_{1}(u_{1},u_{0})^{\frac{1}{2^{n}}}\max(d_{1}(0,u_{0}),d_{1}(0,u_{1}))^{1-\frac{1}{2^{n}}}.$
Proof.
Start with $u_{1}\geq u_{0}$ both smooth Kähler potentials. Let $u_{t}:=(1-t)u_{0}+tu_{1},\ t\in[0,1]$. Differentiating the defining equation of $u^{\beta}_{t}$ (see (16)) we get that $\frac{d}{dt}u^{\beta}_{t}=(u_{1}-u_{0})+\frac{1}{\beta}\Delta_{\omega_{u_{t}^{\beta}}}\frac{d}{dt}u^{\beta}_{t}$. This implies that
$d_{1}(u^{\beta}_{1},u^{\beta}_{0})=I(u^{\beta}_{1})-I(u^{\beta}_{0})=\frac{1}{V}\int_{0}^{1}\int_{X}\frac{d}{dt}u^{\beta}_{t}\omega_{u^{\beta}_{t}}^{n}dt=\frac{1}{V}\int_{0}^{1}\int_{X}(u_{1}-u_{0})\omega_{u^{\beta}_{t}}^{n}dt.$
From [10, Lemma A.2] it follows that
$d_{1}(u^{\beta}_{1},u^{\beta}_{0})\leq C_{n}d_{1}(u_{0},u_{1})^{\frac{1}{2^{n}}}\int_{0}^{1}\max(d_{1}(0,u^{\beta}_{t}),d_{1}(0,u_{0}),d_{1}(0,u_{1}))^{1-\frac{1}{2^{n}}}dt,$
where $C_{n}>0$ is a constant depending only on the dimension.
Using Lemma 3.12, we can write
$d_{1}(0,u_{t}^{\beta})\leq d_{1}(0,u_{t})=-I(u_{t})\leq-(1-t)I(u_{0})-tI(u_{1}),$
where we used the simple fact that $t\mapsto I(u_{t})$ is concave. Thus,
$d_{1}(0,u_{t}^{\beta})\leq\max\{d_{1}(0,u_{0}),d_{1}(0,u_{1})\}.$
As a result, we obtain that
$d_{1}(u^{\beta}_{1},u^{\beta}_{0})\leq C_{n}d_{1}(u_{0},u_{1})^{\frac{1}{2^{n}}}\max(d_{1}(0,u_{0}),d_{1}(0,u_{1}))^{1-\frac{1}{2^{n}}}.$
This can now be extended for $u_{1}\geq u_{0}$, both potentials in $\mathcal{E}^{1}$, using approximation [15].
For general $u_{0},u_{1}\in\mathcal{E}^{1}$, we can write that
$d_{1}(u^{\beta}_{1},P(u_{0},u_{1})^{\beta})$ $\leq C_{n}d_{1}(u_{1},P(u_{0},u_{1}))^{\frac{1}{2^{n}}}\max(d_{1}(0,u_{1}),d_{1}(0,P(u_{0},u_{1})))^{1-\frac{1}{2^{n}}}$
$\leq C^{\prime}_{n}d_{1}(u_{1},u_{0})^{\frac{1}{2^{n}}}\max(d_{1}(0,u_{0}),d_{1}(0,u_{1}))^{1-\frac{1}{2^{n}}},$
where in the last inequality we used that $d_{1}(0,P(u_{0},u_{1}))\leq d_{1}(0,u_{0})+d_{1}(u_{0},P(u_{0},u_{1}))\leq d_{1}(0,u_{0})+d_{1}(u_{0},u_{1})\leq 3\max(d_{1}(0,u_{0}),d_{1}(0,u_{1}))$ .
Therefore, using symmetry, the Pythagorean identity [32] and the fact that $P(u^{\beta}_{0},u^{\beta}_{1})\geq P(u_{0},u_{1})^{\beta}$, we obtain that
$d_{1}(u_{1}^{\beta},u_{0}^{\beta})$ $=d_{1}(u_{1}^{\beta},P(u_{0}^{\beta},u_{1}^{\beta}))+d_{1}(u_{0}^{\beta},P(u_{0}^{\beta},u_{1}^{\beta}))\leq d_{1}(u^{\beta}_{1},P(u_{0},u_{1})^{\beta})+d_{1}(u^{\beta}_{0},P(u_{0},u_{1})^{\beta})$
$\leq C^{\prime\prime}_{n}d_{1}(u_{1},u_{0})^{\frac{1}{2^{n}}}\max(d_{1}(0,u_{0}),d_{1}(0,u_{1}))^{1-\frac{1}{2^{n}}}.$
Using the above result, we obtain that $Ent^{\beta}\{\cdot\}$ and $K^{\beta}\{\cdot\}$ are $d_{1}^{c}$-continuous, the main result of this section. The following argument shows that they are in fact Hölder continuous.
Theorem 5.2.
Let $\{u^{m}_{t}\},\{u_{t}\}\in\mathcal{R}^{1}$ such that $d_{1}^{c}(\{u^{m}_{t}\},\{u_{t}\})\to 0$. Then
$Ent^{\beta}\{u_{t}\}=\lim_{m\to\infty}Ent^{\beta}\{u^{m}_{t}\},\ \ \ K^{\beta}\{u_{t}\}=\lim_{m\to\infty}K^{\beta}\{u^{m}_{t}\}.$ (27)
Proof.
Recall from [42, Lemma 3.2] that for $\{v_{t}\}\in\mathcal{R}^{1}$ there exists $c_{v}\in\mathbb{R}$ with $\sup_{X}v_{t}=c_{v}t$ for $t>0$, and $\{v_{t}\}\mapsto c_{v}$ is $d_{1}^{c}$-continuous. Thus, after possibly replacing $u_{t}$ and $u^{m}_{t}$ by $u_{t}-ct$ and $u^{m}_{t}-ct$ for sufficiently large $c>0$, we may assume that $u_{t},u^{m}_{t}\leq 0$.
Using the previous result and Proposition 3.6, for $l>0$ fixed, we have:
$\Big|\frac{Ent^{\beta}(u_{l})-Ent^{\beta}(u^{m}_{l})}{l}\Big|\leq C_{n}\beta\frac{d_{1}(u_{l},u^{m}_{l})^{\frac{1}{2^{n}}}}{l^{\frac{1}{2^{n}}}}\max\Big(\frac{d_{1}(0,u_{l})}{l},\frac{d_{1}(0,u^{m}_{l})}{l}\Big)^{1-\frac{1}{2^{n}}}+\beta\frac{d_{1}(u_{l},u^{m}_{l})}{l}.$
By Buseman convexity, $\frac{d_{1}(u_{l},u^{m}_{l})}{l}\leq d_{1}^{c}(\{u_{t}\},\{u^{m}_{t}\})\to 0$ for fixed $l>0$, hence:
$\big|Ent^{\beta}\{u_{t}\}-Ent^{\beta}\{u^{m}_{t}\}\big|\leq$ $C_{n}\beta d_{1}^{c}(\{u_{t}\},\{u^{m}_{t}\})^{\frac{1}{2^{n}}}\max(d^{c}_{1}(\{0\},\{u_{t}\}),d^{c}_{1}(\{0\},\{u^{m}_{t}\}))^{1-\frac{1}{2^{n}}}$
$+\beta d_{1}^{c}(\{u_{t}\},\{u^{m}_{t}\}).$
The $d_{1}^{c}$-continuity of $Ent^{\beta}\{\cdot\}$ follows. The $d_{1}^{c}$-continuity of $K^{\beta}\{\cdot\}$ immediately follows from the next lemma. ∎
Lemma 5.3.
Let $u,v\in\mathcal{E}^{1}$ with $u,v\leq 0$. Then
$\big|I_{\operatorname{Ric}\omega}(u)-I_{\operatorname{Ric}\omega}(v)\big|\leq Cd_{1}(u,v)^{\frac{1}{2^{n}}}\max(d_{1}(0,u),d_{1}(0,v))^{1-\frac{1}{2^{n}}}.$
Proof.
Using the cocycle formula of $I_{\operatorname{Ric}\omega}(u)$ we obtain existence of $C(X,\omega)>0$ such that
$C|I_{\operatorname{Ric}\omega}(u)-I_{\operatorname{Ric}\omega}(v)\big|\leq\int_{X}|u-v|\omega^{n}_{\frac{u+v}{4}}=\int_{X}(\max(u,v)-u)+(\max(u,v)-v)\omega^{n}_{\frac{u+v}{4}}.$
Separating the last integral, we estimate the first term. Using [10, Lemma A.2] the same way as in (21) we obtain
$\int_{X}(\max(u,v)-u)$ $\omega^{n}_{\frac{u+v}{4}}\leq Cd_{1}(\max(u,v),u)^{\frac{1}{2^{n}}}\max(J(u),J(\max(u,v)),J((u+v)/4))^{1-\frac{1}{2^{n}}}$
$\leq Cd_{1}(\max(u,v),u)^{\frac{1}{2^{n}}}\max(-I(u),-I(\max(u,v)),-I((u+v)/4))^{1-\frac{1}{2^{n}}}.$
Where in the last line we used that $0\geq\max(u,v),u,(u+v)/4$. Using concavity and monotonicity of $I(\cdot)$, together with $d_{1}(0,u)=-I(u),d_{1}(0,v)=-I(v)$, we arrive at:
$\int_{X}(\max(u,v)-u)\omega^{n}_{\frac{u+v}{4}}$ $\leq Cd_{1}(\max(u,v),u)^{\frac{1}{2^{n}}}\max(d_{1}(0,u),d_{1}(0,v))^{1-\frac{1}{2^{n}}}.$
Similarly, $\int_{X}(\max(u,v)-v)\omega^{n}_{\frac{u+v}{4}}\leq Cd_{1}(\max(u,v),v)^{\frac{1}{2^{n}}}\max(d_{1}(0,u),d_{1}(0,v))^{1-\frac{1}{2^{n}}}.$ After adding the last two estimates, the result follows from [32, Remark 5.6]. ∎
As pointed out in §1, combining Theorem 5.2 with Theorem 4.4, [26, 27] and [61], the proof of our YTD correspondence (Theorem 1.1) is now complete.
6 A transcendental YTD correspondence
We briefly discuss the YTD correspondence for transcendental $(X,\omega)$, not necessarily induced by an ample line bundle. In this case one can still introduce notions of test configurations and K-stability [47, 79].
Definition 6.1.
A Kähler test configuration $\mathcal{T}$ is a pair $\mathcal{T}:=(\mathcal{X},\mathcal{A})$ such that:
$\bullet$ $\mathcal{X}$ is a normal, compact, complex space, $\pi:\mathcal{X}\to\mathbb{P}^{1}$ is a flat, surjective morphism, and there exists a holomorphic $\mathbb{C}^{*}$-action $\rho$ on $\mathcal{X}$ compatible with the standard $\mathbb{C}^{*}$-action on $\mathbb{P}^{1}$.
$\bullet$ There exists a $\mathbb{C}^{*}$-equivariant biholomorphism
$\mathcal{X}\setminus\pi^{-1}(0)\ \cong\ X\times(\mathbb{P}^{1}\setminus\{0\}),$
identifying $\pi$ with projection onto $\mathbb{P}^{1}$.
$\bullet$ $\mathcal{A}\in H^{1,1}_{BC}(\mathcal{X},\mathbb{R})$ is a $\mathbb{C}^{*}$-invariant Bott-Chern cohomology class, and
$\mathcal{A}|_{X\times(\mathbb{P}^{1}\setminus\{0\})}=p_{1}^{*}[\omega],$
where $p_{1}$ is projection to $X$.
$\bullet$ $\mathcal{A}$ is relatively Kähler, i.e. $\mathcal{A}+c\pi^{*}c_{1}(\mathcal{O}_{{\mathbb{P}}^{1}}(1))$ is a Kähler class for some $c>0$.
As in the ample case, it is possible to associate to $\mathcal{T}$ a $C^{1,1}$ geodesic ray $\{u^{\mathcal{T}}_{t}\}$ ([79, Lemma 4.6], [30]), allowing to define $K^{\beta}$-stability in this context in the same way as in the ample case (recall (9)).
A combination of [65, Proposition 6.3.1] and [65, Theorem 5.1.7] implies that finite energy geodesic rays with bounded K energy slope are $d_{1}^{c}$-approximable by rays of Kähler test configurations. As a result, the exact same proof as that of Theorem 1.1 implies the following:
Theorem 6.2.
Let $(X,\omega)$ be a compact Kähler manifold. The following are equivalent:
(i) $X$ admits a unique cscK metric in $\{\omega\}$.
(ii) $X$ is uniformly $K^{\beta}$-stable for some $\beta>0$.
7 A slope formula for the $\beta$-entropy
The rest of this paper is devoted to establishing radial and algebraic formulations of $K^{\beta}$-stability. Let $(X,L)$ be a polarized manifold, with $\omega\in c_{1}(L)$. In this technical section we provide a useful slope formula for our $\beta$-entropy, which is the radial version of (18):
Theorem 7.1.
For any $\beta>1$ and any bounded geodesic ray $\{u_{t}\}$, we have that
$Ent^{\beta}\{u_{t}\}=\sup_{\{v_{t}\}}\left(L^{\beta}(\{v_{t}\},\{u_{t}\})+\beta(I\{v_{t}\}-I\{u_{t}\})\right),$ (28)
where the sup is over all geodesic rays $\{v_{t}\}$ arising from semi-ample test-configurations of $(X,L)$, and
$L^{\beta}(\{v_{t}\},\{u_{t}\}):=\liminf_{t\to\infty}\left(-\frac{1}{t}\log\int_{X}e^{\beta(v_{t}-u_{t})}\frac{\omega^{n}}{V}\right).$ (29)
The boundedness assumption on $\{u_{t}\}$ is naturally satisfied when it comes to applications, since it holds for a large class of geodesic rays (including the ones induced from filtrations; see [76, §9]). That being said, it would be nice to know if this condition is necessary.
The proof of Theorem 7.1 relies on several technical ingredients. The main idea is that, although the ray $\{u_{t}^{\beta}\}$ attains the supremum in (28), it is far from being geodesic. We therefore approximate $\{u_{t}^{\beta}\}$ by a sequence of geodesic rays arising from test configurations.
We begin with a slope formula for $L^{\beta}(\{v_{t}\},\{u_{t}\})$, generalizing the corresponding formula for the Ding energy [5, 42]. As shown later (Theorem 8.1), when both $\{v_{t}\}$ and $\{u_{t}\}$ arise from test configurations, $L^{\beta}(\{v_{t}\},\{u_{t}\})$ can be expressed more precisely as the log canonical threshold of certain divisors.
Proposition 7.2.
Let $\{u_{t}\}$ and $\{v_{t}\}$ be two sublinear finite energy subgeodesic rays. Suppose one of these subgeodesics is a bounded geodesic ray. Then
$L^{\beta}(\{v_{t}\},\{u_{t}\})=\sup\left\{\tau\in\mathbb{R}\ :\int_{0}^{\infty}e^{\tau t}\int_{X}e^{\beta(v_{t}-u_{t})}\omega^{n}dt<\infty\right\}.$
Proof.
We can assume that both sides are finite. Denote the right hand side by $\tau_{0}$.
For any $\tau<L^{\beta}(\{v_{t}\},\{u_{t}\})$, pick $\varepsilon>0$ such that $\tau+\varepsilon<L^{\beta}(\{v_{t}\},\{u_{t}\})$. Then for $t\gg 1$,
$\log\int_{X}e^{\beta(v_{t}-u_{t})}\omega^{n}\geq\tau t+\varepsilon t.$
Thus for $t_{0}$ big enough, $\int_{t_{0}}^{\infty}e^{\tau t}\int_{X}e^{\beta(v_{t}-u_{t})}\omega^{n}dt\leq\int_{t_{0}}^{\infty}e^{-\varepsilon t}dt<\infty.$ So, $L^{\beta}(\{v_{t}\},\{u_{t}\})\leq\tau_{0}$.
For the converse inequality, we rely on the boundedness assumption. In what follows, we suppose that $\{u_{t}\}$ is a bounded ray, the proof in the other case is the same.
Since we are dealing with sublinear subgeodesic rays, we can assume that $t\to u_{t}$ and $t\to v_{t}$ are both $t$-decreasing. Also, [33, Theorem 1] gives some $\gamma>0$ such that
$|u_{t}(x)-u_{t^{\prime}}(x)|\leq\gamma|t-t^{\prime}|,\ \ x\in X$ (30)
Suppose that $\tau<\tau_{0}$. Fixing $t_{0}>1$, we can use the above inequality to start writing
$\int_{t_{0}-1}^{t_{0}}e^{\tau t}$ $\int_{X}e^{\beta(v_{t_{0}}-u_{t_{0}}-\gamma)}\omega^{n}dt\leq\int_{t_{0}-1}^{t_{0}}e^{\tau t}\int_{X}e^{\beta(v_{t_{0}}-u_{t_{0}-1})}\omega^{n}dt$
$\leq\int_{t_{0}-1}^{t_{0}}e^{\tau t}\int_{X}e^{\beta(v_{t}-u_{t})}\omega^{n}dt\leq\int_{0}^{\infty}e^{\tau t}\int_{X}e^{\beta(v_{t}-u_{t})}\omega^{n}dt<C<\infty.$
Carrying out the integration in $t$ on the left hand side we arrive at
$\frac{e^{-\beta\gamma+t_{0}\tau}}{\tau}\big(1-e^{-\tau}\big)\int_{X}e^{\beta(v_{t_{0}}-u_{t_{0}})}\omega^{n}\leq C.$
Taking $\log$ we conclude at
$\log\int_{X}e^{\beta(v_{t_{0}}-u_{t_{0}})}\omega^{n}\geq t_{0}\tau-\beta\gamma+\log\frac{1-e^{-\tau}}{\tau}-\log C.$
Dividing by $t_{0}$ and letting $t_{0}\to\infty$ we arrive at $\liminf_{t_{0}\to\infty}-\frac{1}{t_{0}}\log\int_{X}e^{\beta(v_{t_{0}}-u_{t_{0}})}\omega^{n}\geq\tau.$ So we obtain $L^{\beta}(\{v_{t}\},\{u_{t}\})\geq\tau_{0}$, finishing the proof. ∎
Now we move on to proving Theorem 7.1, which will be carried out in three steps.
Step 1. Replacing $\{u_{t}^{\beta}\}$ with a maximal geodesic ray.
For this step we fix a geodesic ray $\{u_{t}\}\in\mathcal{R}^{1}$ normalized by $\sup u_{t}=0$ (this is possible due to [40, Lemma 3.2]).
For any $\beta>1$, consider the (non-geodesic) ray $\{u_{t}^{\beta}\}$ arising from solving (16). The goal of this step is to find a geodesic replacement for $\{u_{t}^{\beta}\}$, with favorable properties.
By Proposition 4.1, the definition of $Ent^{\beta}\{u_{t}\}$ and (26), it is possible to choose a sequence $t_{j}\to\infty$ such that
$\lim_{j\to\infty}\frac{d_{1}(u_{t_{j}}^{\beta},u_{t_{j}})}{t_{j}}=\lim_{j\to\infty}\frac{Ent^{\beta}(u_{t_{j}})}{\beta t_{j}}=\lim_{j\to\infty}\frac{I(u^{\beta}_{t_{j}})-I(u_{t_{j}})}{t_{j}}=\frac{Ent^{\beta}\{u_{t}\}}{\beta}<\infty.$ (31)
Now let $0,t_{j}]\ni t\mapsto\phi_{j,t,\beta}\in\mathcal{E}^{1}$ be the weak geodesic segment connecting 0 and $u^{\beta}_{t_{j}}$. Due to the convexity of K energy and Corollary 4.5, for $t<t_{j}$ we have
$\frac{K(\phi_{j,t,\beta})}{t}\leq\frac{K(u^{\beta}_{t_{j}})}{t_{j}}\leq\frac{(1+\varepsilon_{\beta})K^{\beta}(u_{t_{j}})+\varepsilon_{\beta}d_{1}(0,u_{t_{j}})+\varepsilon_{\beta}}{t_{j}}.$
So by a standard compactness argument (cf. [36, 10, 37]), we can extract a subsequence, again denoted by $\{\phi_{j,t,\beta}\}$, and obtain a finite energy geodesic ray $\{\phi_{t,\beta}\}$ such that
$\frac{K(\phi_{t,\beta})}{t}\leq(1+\varepsilon_{\beta})K^{\beta}\{u_{t}\}-\varepsilon_{\beta}I\{u_{t}\}<\infty\ \ \textup{ and }\ \ \phi_{j,t,\beta}\xrightarrow{d_{1}}\phi_{t,\beta}\text{ as }j\to\infty.$
Then for any $t>0$, using (31), we obtain a chain of inequalities:
$\begin{split}\frac{I(\phi_{t,\beta})-I(u_{t})}{t}&\leq\frac{d_{1}(\phi_{t,\beta},u_{t})}{t}=\lim_{j\to\infty}\frac{d_{1}(\phi_{j,t,\beta},u_{t})}{t}\leq\lim_{j\to\infty}\frac{d_{1}(u_{t_{j}}^{\beta},u_{t_{j}})}{t_{j}}\\ &=\lim_{j\to\infty}\frac{I(u^{\beta}_{t_{j}})-I(u_{t_{j}})}{t_{j}}=\lim_{j\to\infty}\frac{I(\phi_{j,t,\beta})-I(u_{t})}{t}=\frac{I(\phi_{t,\beta})-I(u_{t})}{t},\end{split}$
where in the second inequality we used the convexity of $t\mapsto d_{1}(\phi_{j,t,\beta},u_{t})$ [11, Proposition 5.1] and in the second line we used the linearity of $t\mapsto I(\phi_{j,t,\beta})-I(u_{t})$.
As a consequence, we obtain equality everywhere:
$\frac{I(\phi_{t,\beta})-I(u_{t})}{t}=\frac{d_{1}(\phi_{t,\beta},u_{t})}{t}=\lim_{j\to\infty}\frac{I(u^{\beta}_{t_{j}})-I(u_{t_{j}})}{t_{j}}\text{ for any }t>0.$ (32)
Therefore, due to (31),
$Ent\{u_{t}\}\geq\lim_{j\to\infty}\frac{Ent^{\beta}(u_{t_{j}})}{t_{j}}=Ent^{\beta}\{u_{t}\}=\beta(I\{\phi_{t,\beta}\}-I\{u_{t}\})=\beta d_{1}^{c}(\{\phi_{t,\beta}\},\{u_{t}\}),$ (33)
and
$K\{\phi_{t,\beta}\}\leq(1+\varepsilon_{\beta})K^{\beta}\{u_{t}\}-\varepsilon_{\beta}I\{u_{t}\}<\infty.$ (34)
[…]
Due to (32) we have that $I(\phi_{t,\beta})-I(u_{t})=I(\phi_{t,\beta})+I(u_{t})-2I(P(u_{t},\phi_{t,\beta}))$. Using the $\mathcal{E}^{1}$ version of [33, Proposition 2.3] we obtain that $P(u_{t},\phi_{t,\beta})=u_{t}$. From this and the fact that $u^{\beta}_{t_{j}}\leq 0$ we obtain that $\phi_{t,\beta}\leq 0$ and
$0\geq\phi_{t,\beta}\geq u_{t}\ \textup{ and }\ \sup_{X}\phi_{t,\beta}=\sup_{X}u_{t}=0.$ (35)
A byproduct of the above construction is the following important theorem. In the setting of algebraic quantization, analogous results have been conjectured by the second author in [95, Conjecture 3.3], but are still open to our knowledge.
Theorem 7.3 (=Theorem 1.2).
Let $\{u_{t}\}\in\mathcal{R}^{1}$. Then
$\lim_{\beta\to\infty}K^{\beta}\{u_{t}\}=K\{u_{t}\}.$
Proof.
Without loss of generality we can assume that $\sup_{X}u_{t}=0,\ t\geq 0$.
When $Ent\{u_{t}\}<\infty$, by the above construction, for any $\beta>1$ the geodesic rays $\{\phi_{t,\beta}\}$ satisfy $d_{1}^{c}(\{\phi_{t,\beta}\},\{u_{t}\})\to 0$ as $\beta\to\infty$ (recall (33). Moreover, due to (34)) we have that
$K\{\phi_{t,\beta}\}\leq(1+\varepsilon_{\beta})K^{\beta}\{u_{t}\}-\varepsilon_{\beta}I\{u_{t}\}\leq(1+\varepsilon_{\beta})K\{u_{t}\}-\varepsilon_{\beta}I\{u_{t}\}.$
Now sending $\beta\to\infty$ and using that $K\{\cdot\}$ is $d_{1}^{c}$-lsc ([25, Proposition 5.9]) we obtain that $\lim_{\beta\to\infty}K^{\beta}\{u_{t}\}=K\{u_{t}\}.$
If $Ent\{u_{t}\}=\infty$, we claim that $\lim_{\beta\to\infty}Ent^{\beta}\{u_{t}\}=\infty$ as well. Indeed, suppose otherwise, then Proposition 3.10 implies that $Ent^{\beta}\{u_{t}\}\leq C$ for some $C>0$ independent of $\beta>0$. So (33) implies that $d_{1}^{c}(\{\phi_{t,\beta}\},\{u_{t}\})\to 0$ as $\beta\to\infty$. Sending $\beta\to\infty$ in (34) and using that $K\{\cdot\}$ is $d_{1}^{c}$-lsc and $K^{\beta}\{u_{t}\}\leq C$, we obtain $K\{u_{t}\}<\infty$, a contradiction. So we conclude. ∎
Step 2. Integrability estimates for $\{\phi_{t,\beta}\}$.
For this step we again fix a geodesic ray $\{u_{t}\}\in\mathcal{R}^{1}$ normalized by $\sup u_{t}=0$. This part is devoted to proving the following integrability condition for the geodesic ray $\{\phi_{t,\beta}\}$ constructed in the previous step.
Proposition 7.4.
For any $\varepsilon>0$ and $\beta>1$,
$\int_{0}^{\infty}e^{-2\varepsilon t}\int_{X}e^{(1-\varepsilon)\beta(\phi_{t,\beta}-u_{t})}\omega^{n}dt<\infty,$
where $\{\phi_{t,\beta}\}$ is the geodesic ray constructed in Step 1.
Recall that, by Proposition 3.14, there exists $A\geq 0$ (depending only on $(X,\omega)$), such that for all $\beta\geq A$ the curve
$t\mapsto u_{t,\beta}:=((1-A/\beta)u_{t})^{\beta}$
is a sublinear finite energy subgeodesic ray. When $\operatorname{Ric}(\omega)\geq 0$, we can simply take $A=0$. Let
$\{v_{t,\beta}\}$ (36)
be the smallest geodesic ray lying above $\{u_{t,\beta}\}$. Such $\{v_{t,\beta}\}$ can be explicitly constructed as the increasing limit of the geodesic segments $0,k]\ni t\mapsto\phi_{k}(t)$ joining 0 and $u_{k,\beta}$ (for details, see the proof of [42, Proposition 3.13]).
Then for any $t>0$, using that K energy is $d_{1}$-lsc and convex we can write
$\begin{split}\frac{K(v_{t,\beta})}{t}\leq\liminf_{k\to\infty}\frac{K(\phi_{k}(t))}{t}\leq\liminf_{k\to\infty}\frac{K(u_{k,\beta})}{k}.\end{split}$
Then by Corollary 4.5 and (the proof of) (26),
$K\{v_{t,\beta}\}<\infty.$
Thus, $\{v_{t,\beta}\}$ must be maximal in the sense of [10], due to [61, Theorem 1.2]. Since $\{v_{t,\beta}\}$ is the smallest possible ray dominating the subgeodesic $\{u_{t,\beta}\}$, in the jargon of [10], $\{v_{t,\beta}\}$ is the maximal geodesic ray associated to the subgeodesic $\{u_{t,\beta}\}$.
Thus the following estimate holds, thanks to [61, Lemma 3.1]:
Lemma 7.5.
For any $\alpha>0$ we have that
$\int_{0}^{\infty}e^{-t}\int_{X}e^{\alpha(v_{t,\beta}-u_{t,\beta})}\omega^{n}dt<\infty.$
Proof.
[61, Lemma 3.1] proves this for geodesic rays, but the same argument works for the subgeodesic ray $\{u_{t,\beta}\}$ (c.f. (15)), since $u_{t,\beta}\leq 0$ due to Lemma 3.1(ii) and $u_{t}\leq 0$ . ∎
Corollary 7.6.
For any $\varepsilon>0$, $\beta>A$ we have that
$\int_{0}^{\infty}e^{-2\varepsilon t}\int_{X}e^{(1-\varepsilon)\beta(v_{t,\beta}-(1-A/\beta)u_{t})}\omega^{n}dt<\infty.$
Proof.
Put for simplicity $Y:=X\times(0,\infty)$ and $d\mu:=\omega^{n}\wedge dt$. By Hölder’s inequality,
[…]
finishing the proof by the previous lemma and the definition of $u_{t,\beta}$. ∎
The next simple comparison principle will play a key role in our argument.
Lemma 7.7.
For any $\delta\in(0,1)$, $\beta>0$ and $u\in\mathcal{E}^{1}$, we have that
$(1-\delta)u^{\beta}\leq((1-\delta)u)^{\frac{\beta}{1-\delta}}-\frac{n(1-\delta)}{\beta}\log(1-\delta).$
Proof.
Consider $(1-\delta)u^{\beta}$, which satisfies
$(\omega+\mathrm{dd^{c}}(1-\delta)u^{\beta})^{n}\geq(1-\delta)^{n}e^{\beta(u^{\beta}-u)}\omega^{n}=e^{\frac{\beta}{1-\delta}((1-\delta)u^{\beta}+\frac{n(1-\delta)}{\beta}\log(1-\delta)-(1-\delta)u)}\omega^{n}.$
On the other hand, $((1-\delta)u)^{\frac{\beta}{1-\delta}}$ satisfies
$\big(\omega+\mathrm{dd^{c}}((1-\delta)u)^{\frac{\beta}{1-\delta}}\big)^{n}=e^{\frac{\beta}{1-\delta}\big(((1-\delta)u)^{\frac{\beta}{1-\delta}}-(1-\delta)u\big)}\omega^{n}.$
So we conclude by the comparison principle [6, Lemma 2.5]. ∎
Let us fix $\delta\in(0,1)$:
$\delta:=\frac{A}{A+\beta}.$
Then $\frac{\beta}{1-\delta}=\beta+A$ and Proposition 3.14 implies that $t\mapsto((1-\delta)u_{t})^{\frac{\beta}{1-\delta}}=((1-\frac{A}{\beta+A})u_{t})^{\beta+A}$ is a subgeodesic ray for any $\beta>0$. By our notation (36), the smallest geodesic ray lying above it was denoted by $\{v_{t,\beta+A}\}$.
We claim that $(1-\delta)\phi_{t,\beta}\leq v_{t,\beta+A}.$ Indeed, recall that $\phi_{t,\beta}$ is constructed as the limit of geodesic segments $\phi_{j,t,\beta}$ joining 0 and $u^{\beta}_{t_{j}}$. Then
$0,t_{j}]\ni t\mapsto(1-\delta)\phi_{j,t,\beta}$
is a subgeodesic segment joining 0 and $(1-\delta)u_{t_{j}}^{\beta}$. Due to Lemma 7.7 and the comparison principle of weak geodesics, this subgeodesic is bounded from above by the geodesic segment joining 0 and $((1-\delta)u_{t_{j}})^{\frac{\beta}{1-\delta}}-\frac{n(1-\delta)}{\beta}\log(1-\delta)$. Sending $j\to\infty$, the constant term $\frac{n(1-\delta)}{\beta}\log(1-\delta)$ will disappear in this limit process, and we obtain the claimed key relation:
$(1-\delta)\phi_{t,\beta}\leq v_{t,\beta+A}.$ (37)
Now, using (37), we obtain the estimate of Proposition 7.4 as follows:
$\int_{0}^{\infty}e^{-2\varepsilon t}\int_{X}e^{(1-\varepsilon)\beta(\phi_{t,\beta}-u_{t})}\omega^{n}dt$ $\leq\int_{0}^{\infty}e^{-2\varepsilon t}\int_{X}e^{(1-\varepsilon)\frac{\beta}{1-\delta}(v_{t,\beta+A}-(1-\delta)u_{t})}\omega^{n}dt$
$=\int_{0}^{\infty}e^{-2\varepsilon t}\int_{X}e^{(1-\varepsilon)(\beta+A)(v_{t,\beta+A}-(1-\frac{A}{A+\beta})u_{t})}\omega^{n}dt<\infty.$
That the last integral is finite follows from Corollary 7.6 with $\beta$ replaced with $\beta+A$. So the proof of Proposition 7.4 is complete.
Corollary 7.8.
For any $\beta>1$ and any bounded geodesic ray $\{u_{t}\}$, there exists a sequence of geodesic rays $\{\psi_{t,k}\}$ such that
$Ent^{\beta}\{u_{t}\}=\lim_{k\to\infty}(L^{\beta}(\{\psi_{t,k}\},\{u_{t}\})+\beta(I\{\psi_{t,k}\}-I\{u_{t}\})).$
Moreover, the sequence $\{\psi_{t,k}\}$ can be chosen so that $Ent\{\psi_{t,k}\}<\infty,$
$\int_{0}^{\infty}e^{-2\varepsilon_{k}t}\int_{X}e^{\frac{\beta}{1-\varepsilon_{k}}(\psi_{t,k}-u_{t})}\omega^{n}dt<\infty\ \textup{ for some }\ \varepsilon_{k}\to 0,$
and
$\lim_{k\to\infty}\beta(I\{\psi_{t,k}\}-I\{u_{t}\})=Ent^{\beta}\{u_{t}\}.$
Proof.
We apply Proposition 7.4 to the parameter $\beta_{\varepsilon}:=\frac{\beta}{(1-\varepsilon)^{2}}$. Then for any $\varepsilon\in(0,1)$,
$\int_{0}^{\infty}e^{-2\varepsilon t}\int_{X}e^{\frac{\beta}{1-\varepsilon}\left(\phi_{t,\beta_{\varepsilon}}-u_{t}\right)}\omega^{n}dt<\infty.$
Due to (35) we obtain that
$\int_{0}^{\infty}e^{-2\varepsilon t}\int_{X}e^{\beta\left(\phi_{t,\beta_{\varepsilon}}-u_{t}\right)}\omega^{n}dt<\infty.$ (38)
Due to (33) and Proposition 3.10 the geodesic ray $\{\phi_{t,\beta_{\varepsilon}}\}$ satisfies
$\beta(I\{\phi_{t,\beta_{\varepsilon}}\}-I\{u_{t}\})=(1-\varepsilon)^{2}Ent^{\frac{\beta}{(1-\varepsilon)^{2}}}\{u_{t}\}\geq(1-\varepsilon)^{2}Ent^{\beta}\{u_{t}\}.$ (39)
Moreover, by Corollary 3.7, $Ent^{\frac{\beta}{(1-\varepsilon)^{2}}}\{u_{t}\}\leq\frac{1}{(1-\varepsilon)^{2}}Ent^{\beta}\{u_{t}\}.$ Theorefore,
$\beta(I\{\phi_{t,\beta_{\varepsilon}}\}-I\{u_{t}\})\to Ent^{\beta}\{u_{t}\}\text{ as }\varepsilon\to 0.$ (40)
Additinally, using (18) we can write
[…]
Taking liminf on both sides and using (39), we obtain that
$\begin{split}Ent^{\beta}\{u_{t}\}\geq\liminf_{t\to\infty}(\frac{-1}{t}\log\int_{X}e^{\beta(\phi_{t,\beta_{\varepsilon}}-u_{t})}\frac{\omega^{n}}{V})+(1-\varepsilon)^{2}Ent^{\beta}\{u_{t}\}.\end{split}$
This implies that
$(1-(1-\varepsilon)^{2})Ent^{\beta}\{u_{t}\}\geq\liminf_{t\to\infty}(\frac{-1}{t}\log\int_{X}e^{\beta(\phi_{t,\beta_{\varepsilon}}-u_{t})}\frac{\omega^{n}}{V})\geq-2\varepsilon,$
where in the last inequality we used Proposition 7.2 and (38). Choosing $\varepsilon_{k}\searrow 0$ and putting
$\psi_{t,k}:=\phi_{t,\beta_{\varepsilon_{k}}},$
by (26) and (29) we get $\lim_{k\to\infty}L^{\beta}(\{\psi_{t,k}\},\{u_{t}\})=0.$ Then we conclude by (40) and (34). ∎
Step 3. Approximation via test configurations.
In this final step we will approximate each $\{\psi_{t,k}\}$ from the above corollary by a sequence of geodesic rays from test configurations and then complete the proof of Theorem 7.1.
Proposition 7.9.
Let $\{u_{t}\}$ be a bounded geodesic ray with $\sup u_{t}=0$. Assume that $\{\psi_{t}\}$ is a geodesic ray such that $Ent\{\psi_{t}\}<\infty$ and
$\int_{0}^{\infty}e^{-2\varepsilon t}\int_{X}e^{\frac{\beta}{1-\varepsilon}(\psi_{t}-u_{t})}\omega^{n}dt<\infty$
for some $\beta>0$ and $\varepsilon>0$. Then there exists a sequence of geodesic rays $\{w_{t,m}\}$ induced from semi-ample test configurations such that
$I\{w_{t,m}\}\to I\{\psi_{t}\}\text{ as }m\to\infty,$
and
$\int_{0}^{\infty}e^{-4\varepsilon t}\int_{X}e^{\beta(w_{t,m}-u_{t})}\omega^{n}dt<\infty\text{ for all }m\gg 1.$
Proof.
During the proof, $\beta>0$ and $\varepsilon>0$ will be fixed. Since $Ent\{\psi_{t}\}<\infty$, [61, Theorem 1.2] is applicable, and applying Demailly’s regularization to $\{\psi_{t}\}$ gives a sequence of geodesic rays $\{w_{t,m}\}$ induced from semi-ample test configurations such that $d_{1}^{c}(\{w_{t,m}\},\{\psi_{t}\})\to 0$, hence
$I\{w_{t,m}\}\to I\{\psi_{t}\}\text{ as }m\to\infty.$
Moreover, for some sufficiently large $m_{0}>0$, we have that (recall (15) and $t=-\log|z|^{2}$)
$\int_{0}^{\infty}e^{-t}\int_{X}e^{(m+m_{0})w_{t,m}-m\psi_{t}}\omega^{n}dt<\infty$
for all $m\gg 1$. We can and will assume that
$m_{0}>\beta.$
Put for simplicity that $Y:=X\times(0,\infty)$ and $d\mu:=\omega^{n}\wedge dt$. Let $\delta\in(0,\varepsilon)$, which will be fixed below. Then we can start writing
[…]
Here $p,q,r>0$ are chosen such that
$\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1,\ \beta p=m+m_{0},\ \beta p(1-\delta)=m,\ \beta q(1-\delta)=\frac{\beta}{1-\varepsilon}.$
These relations are satisfied by choosing
$p=\frac{m+m_{0}}{\beta},\ \delta=\frac{m_{0}}{m+m_{0}},\ q=\frac{m+m_{0}}{(1-\varepsilon)m},\ r=\frac{m+m_{0}}{\varepsilon m+m_{0}-\beta}.$
Therefore, for $m\gg 1$, we have $\delta<\varepsilon$ and
$\delta p=\frac{m_{0}}{\beta}>1,\ 2\varepsilon q=\frac{2\varepsilon}{1-\varepsilon}\frac{m+m_{0}}{m}>2\varepsilon,\ \beta\delta r=\frac{\beta m_{0}}{\varepsilon m+m_{0}-\beta}<\alpha(X,\omega).$
Here $\alpha(X,\omega)$ denotes Tian’s alpha invariant [82]. Then,
$\left(\int_{Y}e^{\beta pw_{t,m}-\beta p(1-\delta)\psi_{t}-\delta pt}d\mu\right)^{1/p}\leq\left(\int_{Y}e^{(m+m_{0})w_{t,m}-m\psi_{t}-t}d\mu\right)^{1/p}<\infty,$
$\left(\int_{Y}e^{\beta q(1-\delta)(\psi_{t}-u_{t})-2\varepsilon qt}d\mu\right)^{1/q}\leq\left(\int_{Y}e^{\frac{\beta}{1-\varepsilon}(\psi_{t}-u_{t})-2\varepsilon t}d\mu\right)^{1/q}<\infty,$
$\left(\int_{Y}e^{-\beta\delta ru_{t}-\varepsilon rt}d\mu\right)^{1/r}\leq\left(C_{\alpha}\int_{Y}e^{-\varepsilon rt}d\mu\right)^{1/r}<\infty.$
So we conclude that $\int_{Y}e^{\beta(w_{t,m}-u_{t})-4\varepsilon t}d\mu<\infty\text{ for all }m\gg 1,$ finishing the proof. ∎
Now we can prove the main result of this section.
Proof of Theorem 7.1.
After possibly adding a linear term, we can assume that $\sup u_{t}=0$ [40, Lemma 3.2]. Due to the definition of $Ent^{\beta}$ (18) it suffices to find a sequence of geodesic rays $\{w_{t,j}\}$ arising from semi-ample test-configurations such that
$Ent^{\beta}\{u_{t}\}=\lim_{j\to\infty}\left(L^{\beta}(\{w_{t,j}\},\{u_{t}\})+\beta(I\{w_{t,j}\}-I\{u_{t}\})\right).$
By Corollary 7.8, we already found a sequence of geodesic rays $\{\psi_{t,k}\}$ with $Ent\{\psi_{t,k}\}<\infty,$
$\int_{0}^{\infty}e^{-2\varepsilon_{k}t}\int_{X}e^{\frac{\beta}{1-\varepsilon_{k}}(\psi_{t,k}-u_{t})}\omega^{n}dt<\infty,\ \textup{ and }\ \lim_{k\to\infty}\beta(I\{\psi_{t,k}\}-I\{u_{t}\})=Ent^{\beta}\{u_{t}\}.$
Additionally, by Proposition 7.9, for each $k>0$, we can find $m_{k}\gg 1$ and a geodesic ray $\{w_{t,m_{k}}\}$ arising from a semi-ample test configuration such that
$\int_{0}^{\infty}e^{-4\varepsilon_{k}t}\int_{X}e^{\beta(w_{t,m_{k}}-u_{t})}\omega^{n}dt<\infty\ \ \textup{and }\ \ |I\{\psi_{t,k}\}-I\{w_{t,m_{k}}\}|\leq 2^{-k}.$
Consequently, we have that
$\beta(I\{w_{t,m_{k}}\}-I\{u_{t}\})\to Ent^{\beta}\{u_{t}\}\text{ as }k\to\infty.$ (41)
Moreover, for some $\delta_{k}\to 0$, (18) gives
$\begin{split}Ent^{\beta}\{u_{t}\}&\geq\liminf_{t\to\infty}\left(\frac{-1}{t}\log\int_{X}e^{\beta(w_{t,m_{k}}-u_{t})}\frac{\omega^{n}}{V}\right)+\beta(I\{w_{t,m_{k}}\}-I\{u_{t}\})\\ &\geq-4\varepsilon_{k}+Ent^{\beta}\{u_{t}\}-\delta_{k},\end{split}$
where we used Proposition 7.2. This implies that $\lim_{k\to\infty}L^{\beta}(\{w_{t,m_{k}}\},\{u_{t}\})=0.$ Coupled with (41), this finishes the proof. ∎
Remark 7.10.
In this section, the polarization assumption for $(X,\omega)$ is only used to invoke Demailly’s approximation when proving Lemma 7.5 and Proposition 7.9. However, as explained in §6, such an approximation procedure also holds for the general Kähler case, thanks to [65]. As a consequence, the transcendental version of Theorem 7.1 holds as well, where the sup in (28) is then over all geodesic rays $\{v_{t}\}$ attached to Kähler test configurations.
8 Algebraic interpretation of $K^{\beta}$-stability
Assume that $(X,L)$ is a polarized manifold. Let $\omega\in c_{1}(L)$, so $V=L^{n}$. Assume that $\{u_{t}\}$ is a geodesic ray arising from a semi-ample test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$. Then we write
$Ent^{\beta}(\mathcal{X},\mathcal{L}):=Ent^{\beta}\{u_{t}\},\ \mathcal{J}(\mathcal{X},\mathcal{L}):=\mathcal{J}_{\operatorname{Ric}(\omega)}\{u_{t}\},\$
$K^{\beta}_{\mathcal{T}}:=K^{\beta}(\mathcal{X},\mathcal{L}):=Ent^{\beta}(\mathcal{X},\mathcal{L})-\mathcal{J}(\mathcal{X},\mathcal{L}).$
It is known from [71, 18] (see also [61, Proposition 2.38]) that
$\mathcal{J}(\mathcal{X},\mathcal{L})=-\frac{\bar{S}\overline{\mathcal{L}}^{n+1}}{(n+1)V}-\frac{K_{X}\cdot\overline{\mathcal{L}}^{n}}{V}.$ (42)
Here the intersection numbers are taken after a trivial compactification $(\overline{\mathcal{X}},\overline{\mathcal{L}})$ of $(\mathcal{X},\mathcal{L})$ over $\infty\in{\mathbb{P}}^{1}$ and then pull back all the line bundles in question to a common normal model.
In what follows we will give a purely algebraic formula for $Ent^{\beta}(\mathcal{X},\mathcal{L})$ as well, in terms of log canonical thresholds and intersection numbers. By Theorem 7.1,
$Ent^{\beta}(\mathcal{X},\mathcal{L})=\sup_{\{v_{t}\}}\left(L^{\beta}(\{v_{t}\},\{u_{t}\})+\beta(I\{v_{t}\}-I\{u_{t}\})\right),$ (43)
where the sup is over all geodesic rays $\{v_{t}\}$ arising from semi-ample test-configurations. Let us therefore take an auxiliary semi-ample test configuration $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$, with $\{v_{t}\}$ being the associated geodesic ray.
We know from [71, 18] that
$\beta(I\{v_{t}\}-I\{u_{t}\})=\frac{\beta((\overline{\mathcal{L}^{\prime}})^{n+1}-\overline{\mathcal{L}}^{n+1})}{(n+1)V}.$ (44)
Therefore, it remains to deal with the term $L^{\beta}(\{v_{t}\},\{u_{t}\})$. We will express it in terms of certain log canonical threshold, as shown in the next result. This recovers Berman’s formula for the Ding invariant in the Fano case [5, (3.19)], by taking $(\mathcal{X}^{\prime},\mathcal{L}^{\prime}):=(\mathcal{X},(1-\frac{1}{\beta})\mathcal{L}+\frac{1}{\beta}L)$.
Theorem 8.1.
Take a common normal model:
$\mathcal{Z}$$\mathcal{X}^{\prime}$$X\times{\mathbb{C}}$$\mathcal{X}$$\scriptstyle{\mu}$$\scriptstyle{\sigma}$$\scriptstyle{\tau}$
There exist $\mathbb{Q}$-Cartier divisors $F$ and $G$ supported on $\mathcal{Z}_{0}$ such that (cf. [79, Propostion 3.10])
$\tau^{*}p_{1}^{*}L-\sigma^{*}\mathcal{L}\sim_{\mathbb{Q}}F,\ \tau^{*}p_{1}^{*}L-\mu^{*}\mathcal{L}^{\prime}\sim_{\mathbb{Q}}G.$
Define a $\mathbb{Q}$-Weil divisor $\mathcal{D}$ supported in $\mathcal{Z}_{0}$ by
$\mathcal{D}:=\beta(F-G)-K_{\mathcal{Z}}+\tau^{*}K_{X\times{\mathbb{C}}}.$ (45)
Then
$L^{\beta}(\{v_{t}\},\{u_{t}\})=\mathrm{lct}_{(\mathcal{Z},\mathcal{D})}(\mathcal{Z}_{0})-1.$ (46)
(See [17, §1.5] for the definition of log canonical threshold (lct) for a pair.)
Moreover, when $\mathcal{Z}$ is smooth with simple normal crossing (snc) central fiber, we can write
$\mathcal{Z}_{0}=\sum_{i}a_{i}E_{i},\ \ F=\sum_{i}b_{i}E_{i},\ \ G=\sum_{i}c_{i}E_{i},\ \ K_{\mathcal{Z}/X\times{\mathbb{C}}}:=K_{\mathcal{Z}}-\tau^{*}K_{X\times{\mathbb{C}}}=\sum_{i}d_{i}E_{i}.$
In this case, we can write more explicitly that
$L^{\beta}(\{v_{t}\},\{u_{t}\})=\min_{i}\frac{(d_{i}+1-a_{i})+\beta(c_{i}-b_{i})}{a_{i}}.$ (47)
Proof.
Put $U_{z}:=u_{-\log|z|^{2}}$ and $V_{z}:=v_{-\log|z|^{2}}$. Applying the change of variables $t=-\log|z|^{2}$ in the statement of Proposition 7.2, we obtain
$L^{\beta}(\{v_{t}\},\{u_{t}\})=\sup\left\{\tau\in\mathbb{R}:\int_{X\times\Delta^{*}}e^{\beta(V_{z}-U_{z})-(1+\tau)\log|z|^{2}}\omega^{n}\wedge idz\wedge d\bar{z}<\infty\right\}.$
The result then simply follows from expanding all the terms appearing in the above integral.
Indeed, when $\mathcal{Z}$ is smooth with snc central fiber, the complexified geodesic rays $z\mapsto U_{z},\ V_{z}$ when pulled back to $\mathcal{Z}$ have only divisorial singularities along $F$ and $G$, respectively (see [79, Lemma 4.6] and [30, §3]). Then the integrability of $e^{\beta(V_{z}-U_{z})-(1+\tau)\log|z|^{2}}$ is equivalent to requiring that the coefficients of the divisor
$\beta(F-G)-K_{\mathcal{Z}/X\times\mathbb{C}}+(1+\tau)\mathcal{Z}_{0}=\sum_{i}\big(\beta(b_{i}-c_{i})+a_{i}(1+\tau)-d_{i}\big)E_{i}$
are all strictly less than $1$, giving (47).
Lastly, by the pull back formula of pairs (see [17, §1.5]), for a general normal model $\mathcal{Z}$ dominating $\mathcal{X}$, $X\times{\mathbb{C}}$ and $\mathcal{X}^{\prime}$, the quantity $L^{\beta}(\{v_{t}\},\{u_{t}\})+1$ can then be expressed as the lct of the pair $(\mathcal{Z},\ \beta(F-G)-K_{\mathcal{Z}/X\times\mathbb{C}})$ with respect to $\mathcal{Z}_{0}$, giving (46). So we conclude. ∎
Let $\mathcal{T}=(\mathcal{X},\mathcal{L})$ be a semi-ample test configuration. Let $Div_{\mathcal{X}}$ be the set of $\mathbb{Q}$-Cartier divisors $G$ contained in the central fiber $\mathcal{Z}_{0}$ of a $\mathbb{C}^{*}$-equivariant smooth model $\pi:\mathcal{Z}\to\mathcal{X}$, with $\pi$ bijective on $\mathcal{X}\setminus\mathcal{X}_{0}$, and $\mathcal{L}_{G}:=\pi^{*}p_{1}^{*}L-G$ relatively semi-ample on $\mathcal{Z}$.
Thus defined, $(\mathcal{Z},\mathcal{L}_{G})$ is also a semi-ample test configurations, and we define the corresponding divisor $\mathcal{D}_{G}$ as in (45):
$\mathcal{D}_{G}:=\beta(F-G)-K_{\mathcal{Z}}+\tau^{*}K_{X\times{\mathbb{C}}}.$ (48)
Now we plug (46) and (44) into (43). Together with (42), this gives a divisorial threshold interpretation for $K^{\beta}_{\mathcal{T}}=Ent^{\beta}(\mathcal{X},\mathcal{L})-\mathcal{J}(\mathcal{X},\mathcal{L})$:
Corollary 8.2.
Let $\mathcal{T}:=(\mathcal{X},\mathcal{L})$ be a semi-ample test configuration. With the notation introduced above
$K^{\beta}_{\mathcal{T}}=\sup_{G\in Div_{\mathcal{X}}}\Big(\mathrm{lct}_{(\mathcal{\mathcal{Z}},\mathcal{D}_{G})}(\mathcal{Z}_{0})-1+\frac{\beta\overline{\mathcal{L}_{G}}^{n+1}-\beta\overline{\mathcal{L}}^{n+1}}{(n+1)V}\Big)+\frac{\bar{S}\overline{\mathcal{L}}^{n+1}}{(n+1)V}+\frac{K_{X}\cdot\overline{\mathcal{L}}^{n}}{V}.$ (49)
Being the supremum of divisorial data, the above expression is reminiscent of other known divisorial threshold type expressions from the Fano case, like the delta and alpha invariants.
It might be a bit too optimistic to expect that the supremum in (49) is achieved by some $G\in Div_{\mathcal{X}}$. But using a different argument, an intersection type formula for $K^{\beta}_{\mathcal{T}}$ can still be obtained, as will be proved in a sequel paper [43]:
$K^{\beta}_{\mathcal{T}}=\frac{\beta\langle\overline{\mathcal{L}_{\beta}}^{n+1}\rangle-\beta\overline{\mathcal{L}}^{n+1}}{(n+1)V}+\frac{\bar{S}\overline{\mathcal{L}}^{n+1}}{(n+1)V}+\frac{K_{X}\cdot\overline{\mathcal{L}}^{n}}{V},$ (50)
where $\langle\overline{\mathcal{L}_{\beta}}^{n+1}\rangle$ is the top movable intersection product of the (relatively) big line bundle $\overline{\mathcal{L}_{\beta}}:=\overline{\mathcal{L}}+\frac{1}{\beta}\big(K_{\overline{\mathcal{X}}}-\pi^{*}K_{X\times\mathbb{P}^{1}}-\mathcal{X}_{0,red}+\mathcal{X}_{0}\big)$.
$K^{\beta}$-stability vs. Ding stability.
When $(X,L)=(X,-K_{X})$ is Fano, our $K^{\beta}(\mathcal{X},\mathcal{L})=K^{\beta}_{\mathcal{T}}$ is directly related to $\mathrm{Ding}(\mathcal{X},\mathcal{L})$, the Ding invariant introduced in [5, 10]. Indeed, choosing $G:=(1-\frac{1}{\beta})F$ in (48) (i.e., taking $\mathcal{L}_{G}:=(1-\frac{1}{\beta})\mathcal{L}+\frac{1}{\beta}(-K_{X})$), one then has (see [90, (2.18)]) $\mathrm{Ding}(\mathcal{X},\mathcal{L})=\mathrm{lct}_{(\mathcal{\mathcal{Z}},\mathcal{D}_{G})}(\mathcal{Z}_{0})-1-\frac{\overline{\mathcal{L}}^{n+1}}{(n+1)V}.$ Thus, (49) implies that (using $\bar{S}=n$ here)
$\begin{split}K^{\beta}(\mathcal{X},\mathcal{L})&\geq\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{\beta(\overline{\mathcal{L}_{G}}^{n+1}-\overline{\mathcal{L}}^{n+1})}{(n+1)V}+\frac{(\overline{\mathcal{L}}-(-K_{X}))\cdot\overline{\mathcal{L}}^{n}}{V}\\ &\geq\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{\beta}{V}(\overline{\mathcal{L}_{G}}-\overline{\mathcal{L}})\cdot\overline{\mathcal{L}_{G}}^{n}+\frac{(\overline{\mathcal{L}}-(-K_{X}))\cdot\overline{\mathcal{L}}^{n}}{V}\\ &=\mathrm{Ding}(\mathcal{X},\mathcal{L})+\frac{1}{V}(-K_{X}-\overline{\mathcal{L}})\cdot(\overline{\mathcal{L}_{G}}^{n}-\overline{\mathcal{L}}^{n}),\end{split}$
where in the second line we used the Hodge index type estimate [90, Lemma 1.74] (cf. also (2.4) therein) and the last line follows from the relation $\beta(\overline{\mathcal{L}_{G}}-\overline{\mathcal{L}})=-K_{X}$-$\overline{\mathcal{L}}$.
We further argue that the error term $\frac{1}{V}(-K_{X}-\overline{\mathcal{L}})\cdot(\overline{\mathcal{L}_{G}}^{n}-\overline{\mathcal{L}}^{n})$ appearing above conveniently controlled from below when $\beta\gg 1$. Indeed, up to adding some multiple $\mathcal{X}_{0}$ to $\mathcal{L}$, we can assume that $(-K_{X}-\overline{\mathcal{L}})$, when pulled back to a common normal model $\mathcal{Z}$, is an effective $\mathbb{Q}$-divisor supported in $\mathcal{Z}_{0}$ and it does not contain any multiple of $\mathcal{Z}_{0}$ (so that $\overline{\mathcal{L}}\cdot(-K_{X})^{n}=0$). Then using that $\overline{\mathcal{L}_{G}}-(1-\frac{1}{\beta})\overline{\mathcal{L}}$ is nef, we deduce that
$\begin{split}\frac{1}{V}(-K_{X}-\overline{\mathcal{L}})\cdot(\overline{\mathcal{L}_{G}}^{n}-\overline{\mathcal{L}}^{n})\geq-\frac{(1-(1-\beta^{-1})^{n})}{V}(-K_{X}-\overline{\mathcal{L}})\cdot\overline{\mathcal{L}}^{n}=-\varepsilon_{\beta}\mathbf{I}(\mathcal{X},\mathcal{L}),\end{split}$
where $\varepsilon_{\beta}:=(1-(1-\beta^{-1})^{n})$ and $\mathbf{I}(\mathcal{X},\mathcal{L})$ denotes the $\mathbf{I}$-norm of $(\mathcal{X},\mathcal{L})$ (see [90, (2.2)]). Using the relation $\mathbf{I}\leq(n+1)\mathbf{J}$ [90, Proposition 2.9], we finally arrive at
$K^{\beta}(\mathcal{X},\mathcal{L})\geq\mathrm{Ding}(\mathcal{X},\mathcal{L})-\varepsilon^{\prime}_{\beta}\mathbf{J}(\mathcal{X},\mathcal{L}).$
Consequently, uniform Ding stability (meaning that $\mathrm{Ding}(\mathcal{X},\mathcal{L})\geq\gamma\mathbf{J}(\mathcal{X},\mathcal{L})$ for some uniform $\gamma>0$) implies uniform $K^{\beta}$-stability for $\beta\gg 1$. To show the reverse implication, one can use that uniform $K^{\beta}$-stability implies uniform $K$-stability (as $K\geq K^{\beta}$) and then invoke the deep fact that the latter in turn implies uniform Ding stability [10, 53] (which relies on the MMP techniques of Li–Xu [62]). Therefore, the above purely algebraic consideration shows that stability notions defined using $K$, $K^{\beta}$ and $\mathrm{Ding}$ are all equivalent for Fano manifolds.
9 Connections with the non-Archimedean point of view
We briefly point out the relationship of our work with the non-Archimedean point of view and give an alternative argument of a recent remarkable result due to Boucksom–Jonsson [23]. For background and terminology on non-Archimedean geometry we refer to [17].
Assume that $(X,L)$ is a polarized manifold with $\omega\in c_{1}(L)$. Let us first give a non-Archimedean interpretation for the radial $\beta$-entropy. In light of Theorem 7.1, it suffices to give a non-Archimedean formula for $L^{\beta}(\{v_{t}\},\{u_{t}\})$. We achieve this by adapting the treatment of the Ding energy from [10, Theorem 5.4].
Proposition 9.1.
Let $\{u_{t}\}$ be a bounded geodesic ray, and let $\{v_{t}\}$ be a geodesic ray arising from a semi-ample test configuration. Then for any $\beta>0$,
$L^{\beta}(\{v_{t}\},\{u_{t}\})=\inf_{w\in X_{\mathbb{Q}}^{div}}(A_{X}(w)+\beta(u^{an}(w)-v^{an}(w))).$
Here $u^{an}$ and $v^{an}$ are non-Archimedean potentials associated with $\{u_{t}\}$ and $\{v_{t}\}$ in the sense of [10, Definition 4.2] (see also [41, Proposition 3.1]). Moreover, $X^{div}_{\mathbb{Q}}$ denotes the $\mathbb{Q}$-valued divisorial valuations on $X$ (including the trivial one).
Proof.
Since both $\{u_{t}\}$ and $\{v_{t}\}$ are bounded rays, $L^{\beta}(\{v_{t}\},\{u_{t}\})$ is a finite number. Moreover, since the statement we aim to show is translation invariant, we can assume that $u_{t}\leq 0$ and $v_{t}\leq 0$ so that $U_{z}:=u_{-\log|z|^{2}}$ and $V_{z}:=v_{-\log|z|^{2}}$ are qpsh functions on $X\times\Delta$. By Proposition 7.2 and using $t=-\log|z|^{2}$ we have that
$L^{\beta}(\{v_{t}\},\{u_{t}\})=\sup\left\{\tau\in\mathbb{R}:\int_{X\times\Delta}e^{\beta(V_{z}-U_{z})-(1+\tau)\log|z|^{2}}\omega^{n}\wedge idz\wedge d\bar{z}<\infty\right\}.$
Up to further subtracting $Ct$ from $\{v_{t}\}$ we can assume that $e^{\beta(V_{z}-U_{z})}\in L^{1}(X\times\Delta)$. Put for simplicity $\mu:=e^{\beta(V_{z}-U_{z})}\omega^{n}\wedge idz\wedge d\bar{z}$. Then, in terms of the complex singularity exponent (see [42, §2.2]), we can write
$L^{\beta}(\{v_{t}\},\{u_{t}\})+1=c_{\mu}[\log|z|^{2}]:=\sup\left\{\lambda>0:\int_{X\times\Delta}e^{-\lambda\log|z|^{2}}d\mu<\infty\right\}.$
Noting that $V_{z}$ has algebraic singularities (recall (14)) and that $V_{z}$, $U_{z}$ and $\log|z|^{2}$ are all $S^{1}$-invariant and bounded away from $X\times\{0\}$, we can apply the valuative criterion [10, Theorem B.7] in the same way as in the short proof of [42, Theorem 2.3] to deduce that
$c_{\mu}[\log|z|^{2}]=\inf_{E}\frac{A_{X\times\Delta}(E)+\beta(\nu(V_{z},E)-\nu(U_{z},E))}{\nu(\log|z|^{2},E)},$
[…]
$\nu(\phi,E):=\inf_{x\in E}\sup\{c>0:\phi(y)-c\log|y-x|^{2}\text{ is bounded above near }x\}.$
(Our convention of the Lelong number differs from the one used in [10] by a factor of $2$).
Consequently, $L^{\beta}(\{v_{t}\},\{u_{t}\})$ is equal to the largest $\tau\in\mathbb{R}$ such that
$A_{X\times\Delta}(E)+\beta\nu(V,E)-\beta\nu(U,E)\geq(1+\tau)\nu(\log|z|^{2},E)$ (51)
[…]
$b_{E}:=\nu(\log|z|^{2},E).$
By [17, Theorem 4.6], all divisorial valuations $w\in X^{div}_{\mathbb{Q}}$ (including the trivial valuation) are in one-to-one correspondence with $\frac{1}{b_{E}}\mathrm{ord}_{E}$, via the Gauss map/extension. In fact, the Gauss extension $G(w)$ of any $w\in X^{div}_{\mathbb{Q}}$ is of the form $G(w)=\frac{1}{b_{E}}\mathrm{ord}_{E}$ for some $E$ over $X\times\{0\}$ and we have the relation $A_{X}(w)=\frac{1}{b_{E}}A_{X\times\Delta}(E)-1$ (by [21, Theorem A.10]). Moreover, we have that (cf. [10, Definition 4.2])
$u^{an}(w)=-\frac{1}{b_{E}}\nu(U,E)\text{ and }v^{an}(w)=-\frac{1}{b_{E}}\nu(V,E).$
Revisiting (51), we obtain that $L^{\beta}(\{v_{t}\},\{u_{t}\})$ is equal to the largest $\tau\in\mathbb{R}$ such that
$A_{X}(w)+\beta(u^{an}(w)-v^{an}(w))\geq\tau\text{ for all }w\in X^{div}_{\mathbb{Q}}.$
Therefore, $L^{\beta}(\{v_{t}\},\{u_{t}\})=\inf_{w\in X^{div}_{\mathbb{Q}}}(A_{X}(w)+\beta(u^{an}(w)-v^{an}(w))),$ finishing the proof. ∎
Corollary 9.2.
Let $\{u_{t}\},\{v_{t}\}$ be geodesic rays induced by a model filtration and a semi ample test configuration respectively. Then for any $\beta>0$:
$L^{\beta}(\{v_{t}\},\{u_{t}\})=L^{NA}(\beta({u}^{an}-v^{an})),$ (52)
where $L^{NA}(f):=\inf_{w\in X^{div}}(A_{X}(w)+f(w))$ is the natural non-Archimedean functional from [20, §2.6], defined for continuous non-Archimedean functions.
Proof.
Since model filtrations are linearly bounded, the associated ray $\{u_{t}\}$ must be bounded [76], hence the previous result is applicable. Both $v^{an}$ and $u^{an}$ are continuous on $X^{an}$ by the sentence below [61, formula (28)], hence (52) holds, by the definition of $L^{NA}$. ∎
Corollary 9.3.
Let $\{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}=\mathcal{E}^{1,NA}$ be a maximal geodesic ray. Then for any $\beta>1$,
$Ent^{\beta}\{u_{t}\}\leq Ent^{NA}(\mathrm{MA}(u^{an})).$ (53)
Here $\mathrm{MA}(u^{an})$ denotes the non-Archimedean Monge-Ampère measure of $u^{an}$ and $Ent^{NA}$ denotes the non-Archimedean entropy (see [17, §6.7 and Definition 7.17]).
Proof.
Due to Theorem 5.2, [61, Proposition 6.3] and [75, Theorem 4.4.1], we can approximate both sides of (53) using a sequence of geodesic rays arising from model filtrations in the sense of [61, Definition 2.7].
So in what follows we simply assume that $\{u_{t}\}$ itself a geodesic ray arising from a model filtration. Pick any geodesic ray $\{v_{t}\}$ arising from a semi-ample test configuration. Corollary 9.2 yields $L^{\beta}(\{v_{t}\},\{u_{t}\})=L^{NA}(\beta({u}^{an}-v^{an})).$ Applying [20, Proposition 2.8] and [22, (4.5)],
[…]
Taking supremum over all $\{v_{t}\}$, we conclude by Theorem 7.1. ∎
Consequently, we obtain an alternative proof of a recent result of Boucksom–Jonsson [23]:
Corollary 9.4.
For any maximal geodesic ray $\{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}=\mathcal{E}^{1,NA}$ one has
$K\{u_{t}\}=K^{NA}(u^{an}).$ (54)
Proof.
Letting $\beta\to\infty$ in (53), we obtain from Theorem 1.2 that $K\{u_{t}\}\leq K^{NA}(u^{an}).$ Combining with the reverse inequality [61, Theorem 1.2], we conclude the proof. ∎
$K^{\beta}$-stability vs. $\widehat{K}-$stability.
Finally, we show directly that uniform $K^{\beta}$-stability (for $\beta$ sufficiently large) is equivalent with the notion of uniform $\widehat{K}-$stability introduced in [61, 23], without involving cscK metrics.
According to [61, 23] and (54), $(X,L)$ being uniformly $\widehat{K}-$stable means that there exists $\gamma>0$ such that $K\{u_{t}\}\geq\gamma J\{u_{t}\}$ holds for all maximal ray $\{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}}=\mathcal{E}^{1,NA}.$
Let us assume that $(X,L)$ is uniformly $K^{\beta}$-stable. By Theorem 5.2, this implies that $K^{\beta}\{u_{t}\}\geq\gamma J\{u_{t}\}$ holds for all maximal ray $\{u_{t}\}\in\mathcal{R}^{1}_{\mathcal{I}},$ as $\{u_{t}\}$ can be $d_{1}^{c}$-approximated by geodesic rays of ample test configurations [10, 40]. This further implies uniformly $\widehat{K}-$stability, since $K\geq K^{\beta}$ (recall (19)).
Conversely, assume that $(X,L)$ is uniformly $\widehat{K}-$stable. Let $\{u_{t}\}$ be a geodesic ray of an ample test configuration. Without loss of generality, one can assume the normalization $\sup_{X}u_{t}=0,\ t\geq 0$. Consider the maximal geodesic rays $\{\phi_{t,\beta}\}\in\mathcal{R}^{1}_{\mathcal{I}}=\mathcal{E}^{1,NA}$ constructed under (31) for $\beta>1$. Then $K\{\phi_{t,\beta}\}\geq\gamma J\{\phi_{t,\beta}\}$, by uniform $\widehat{K}-$stability. Then the normalization $\sup_{X}u_{t}=0$, (24), and (34) further imply that
$(1+\varepsilon_{\beta})K^{\beta}\{u_{t}\}+\varepsilon_{\beta}J\{u_{t}\}\geq K\{\phi_{t,\beta}\}\geq\gamma J\{\phi_{t,\beta}\}.$
Using (35), (24), (33) and [36, Proposition 2.5], we also have that
$|J\{u_{t}\}-J\{\phi_{t,\beta}\}|=|I\{\phi_{t,\beta}\}-I\{u_{t}\}|=\frac{1}{\beta}Ent^{\beta}\{u_{t}\}\leq\frac{1}{\beta}(K^{\beta}\{u_{t}\}+CJ\{u_{t}\}).$
Putting the last two inequalities together, for any $\varepsilon>0$ there exists $\beta_{\varepsilon}>1$ such that
$K^{\beta}\{u_{t}\}\geq(\gamma-\varepsilon)J\{u_{t}\},\ \ \beta\geq\beta_{\varepsilon},$
showing that $(X,L)$ is uniformly $K^{\beta}$-stable for all sufficiently large $\beta$.
Even though the above reasoning does not go through the existence of cscK metrics, it does utilize (54) and the analytic theory of geodesic rays explored in [37]. It would be desirable to find a completely algebraic proof of the above equivalence.